### v2

# Document-Level In-Context Few-Shot Relation Extraction via Pre-Trained Language Models

Abstract
Relation extraction aims at inferring structured human knowledge from textual documents. State-of-the-art methods based on language models commonly have two limitations: (1) they require named entities to be either given as input or infer them, which introduces additional noise, and (2) they require human annotations of documents. As a remedy, we present a novel framework for document-level in-context few-shot relation extraction via pre-trained language models. We achieve crucial benefits in that we eliminate the need for both named entity recognition and human annotation of documents. Unlike existing methods based on fine-tuning, our framework is flexible in that it can be easily updated for a new set of relations without re-training. We evaluate our framework using DocRED, the largest publicly available dataset for document-level relation extraction, and demonstrate that our framework achieves state-of-the-art performance. Finally, we show that our framework actually performs much better than the original labels from the development set of DocRED. To the best of our knowledge, we are the first to reformulate the document-level relation extraction task as a tailored in-context few-shot learning paradigm.
LLM, relation extraction, in-context learning
[…]
*supplements
1 Introduction
Relational facts are widely used to represent human knowledge (Grishman, 2019; Han et al., 2020; Weikum et al., 2021). With the explosion of the web, relational facts have become broadly available through large knowledge bases (KBs) (Auer et al., 2007; Suchanek et al., 2007; Bollacker et al., 2008; Vrandečić and Krötzsch, 2014) and thereby support many downstream tasks. Examples are commonsense reasoning (Lin et al., 2019; Liu et al., 2021a), question answering (Luo et al., 2018; Das et al., 2022; Wang et al., 2022), fact checking (Huynh and Papotti, 2019; Vedula and Parthasarathy, 2021), and product recommendations (Wang et al., 2018; Zhou et al., 2020a, b). However, relational facts are not readily available in structured form but are commonly embedded in unstructured texts. To this end, methods are needed for relation extraction from text.
State-of-the-art methods for relation extraction leverage pre-trained language models (LMs) and fine-tune them using human-annotated documents. These works can be loosely grouped into two streams. One stream requires named entities (e. g., by specifying the entities of interest with a special token) as input (Wang et al., 2019; Xu et al., 2021; Zhang et al., 2021; Zhou et al., 2021; Tan et al., 2022; Wang Xu and Zhao, 2022; Xiao et al., 2022; Hu et al., 2023; Xu et al., 2023). Another stream avoids the use of named entities as input and, instead, learns the detection of named entities through a tailored training objective (Cabot and Navigli, 2021; Eberts and Ulges, 2021; Giorgi et al., 2022; Lu et al., 2022b; Zhang et al., 2023b).
Yet, state-of-the-art methods based on LMs have three main drawbacks that limit their applicability for relation extraction in practice. (1) State-of-the-art methods for relation extraction typically require the named entities to be either given as input or to be inferred via a customized training objective. This introduces additional noise in the relation extraction pipeline and thus hinders the downstream performance. (2) State-of-the-art methods for relation extraction need large amounts of human-annotated documents for training. However, human annotation is costly. (3) State-of-the-art methods are based on LMs that are fine-tuned. As a result, whenever a new type of relation is added to the knowledge base or whenever a better LM is adopted, the entire training process must be repeated. This introduces a huge computational overhead.
There are some recent efforts that use the reasoning abilities of LMs via in-context learning for relation extraction (Wadhwa et al., 2023; Wan et al., 2023). However, these are designed for sentence-level relation extraction, meaning for a small set of relation types. Hence, due to high computational costs, their scalability to documents is limited (see Table 1). Here, we introduce a novel method to leverage in-context learning for document-level relation extraction.
Our REPLM framework: We introduce a novel framework called REPLM for document-level in-context few-shot relation extraction via pre-trained language models. Our framework leverages the generalization capabilities of pre-trained LMs by reformulating the relation extraction task as a tailored in-context few-shot learning paradigm. Specifically, for a given document, we retrieve sets of the most relevant in-context examples of a corresponding relation and aggregate the outputs in a probabilistic framework.
Contributions: [^1] Codes are available at https://github.com/oezyurty/REPLM .
  1. 1.
    We present a novel framework called REPLM for in-context few-shot relation extraction via pre-trained LMs. To the best of our knowledge, we are the first to reformulate the document-level relation extraction task as a tailored in-context few-shot learning paradigm.
  2. 2.
    Our REPLM framework has key advantages for practice: it eliminates the noise from named entity recognition, it circumvents the need for human annotations, and it is flexible in that it is directly applicable to new relations without re-training.
  3. 3.
    We show that our REPLM framework achieves state-of-the-art performance.
2 Related Work
In-context few-shot learning of LMs: LMs have achieved superior performance in many downstream tasks (Beltagy et al., 2019; Devlin et al., 2019; Liu et al., 2019; Radford et al., 2019; Brown et al., 2020; Lewis et al., 2020; Raffel et al., 2020; Wang and Komatsuzaki, 2021; Wei et al., 2022a, b; OpenAI, 2023; Wang et al., 2023; Zhang et al., 2023c). Due to the large computational cost of fine-tuning an LM, Brown et al. (2020) proposed in-context few-shot learning paradigm to teach an LM a new task at inference time. We provide an overview of applications in Appendix A. However, we are not aware of any earlier work that leveraged the in-context few-shot learning paradigm for document-level relation extraction.
Early research on relation extraction: Early works extracted relations from text via pattern extraction methods (Carlson et al., 2010; Nakashole et al., 2012; Jiang et al., 2017; Pawar et al., 2017; Weikum et al., 2021) and via statistical methods (Sarawagi and Cohen, 2004; Zhang et al., 2006a, b; Jiang and Zhai, 2007; Nguyen et al., 2007; Wang, 2008; Yu and Lam, 2010; Wang et al., 2014; Lin et al., 2015). However, the above methods have only a limited modeling capacity, as compared to neural networks (Zeng et al., 2014; Miwa and Bansal, 2016; Zhou et al., 2016; Adel and Schütze, 2017; Katiyar and Cardie, 2017; Zheng et al., 2017; Han et al., 2020). As shown later, LM-based methods better capture the complex interactions between named entities to classify the relation. A detailed review is in the Appendix A.
Relation extraction via LM: State-of-the-art methods for relation extraction are based on fine-tuning pre-trained LMs. Specifically, these methods use pre-trained LMs such as BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), and SciBERT (Beltagy et al., 2019) and fine-tuned them for relation extraction. For instance, Wang et al. (2019) fine-tuned BERT to classify the relation between each named entity pair in a sentence. There have been various follow-up works to improve performance by learning complex dependency between named entities (Wang and Lu, 2020; Paolini et al., 2021; Xu et al., 2021; Zhang et al., 2021; Zhou et al., 2021; Tan et al., 2022; Wang Xu and Zhao, 2022; Xiao et al., 2022; Hu et al., 2023; Xu et al., 2023; Zhang et al., 2023a). A detailed review is in Appendix A. However, all these works require the named entities to be annotated and provided as input at both training and test time.
Some works relax the requirement of given named entities to facilitate processing the raw documents at test time. As a remedy, these works jointly learn to extract named entities and relations. Examples are SpERT (Eberts and Ulges, 2020), JEREX (Eberts and Ulges, 2021), Seq2Rel (Giorgi et al., 2022), UIE (Lu et al., 2022b), and TaG (Zhang et al., 2023b). Their drawback is that multi-step pipelines with named entities recognition propagate the errors to relation extraction (Cabot and Navigli, 2021). Motivated by this, Cabot and Navigli (2021) developed REBEL, an auto-regressive model based on BART (Lewis et al., 2020), which is fine-tuned to output relations as sequences of texts. To the best of our knowledge, REBEL is the only LM-based method without the need for a named entity recognition pipeline and thus represents our main baseline.
Still, the above methods have salient limitations: (1) they (with the exception of REBEL) require named entities to be given or infer them, which is a source of noise; (2) they require large amounts of human annotations; and (3) they require re-training to handle new relations.
Table 1: Comparison of relevant relation extraction methods.
Method Scope Pre-trained LM Open source No need for fine-tuning. No need for named entities.
Wan et al. (2023) Sentence OpenAI’s GPT
Wadhwa et al. (2023) Sentence OpenAI’s GPT
REBEL (Cabot et al., 2021) Document BART-large
REPLM (ours) Document GPT-J
In-context learning for sentence-level relation extraction:. There are two recent works that leverage in-context learning for sentencel-level relation extraction (see Table 1).Wan et al. (2023) requires named entities to be provided for each sentence and generates $k$ different chain-of-thought (CoT) reasonings (Wei et al., 2022b) from GPT-3 (Brown et al., 2020) as in-context examples for each named entity pairs to classify the relation between them. Similarly, Wadhwa et al. (2023) requires CoT reasonings from GPT-3 for the entire training corpus and fine-tunes Flan-T5 (Chung et al., 2022) based on the generated CoT outputs. However, these works are only applicable to a small set of sentences and relation types due to their high computational cost resulting from commercial architectures. Because of that, these methods work only at the sentence-level relation extraction (i. e., they are not scalable to document-level).
Research gap: To the best of our knowledge, no work has adapted the in-context few-shot learning paradigm for document-level relation extraction. This presents our novelty and offers direct benefits in practice (i.e., no need for named entity input, no need for human annotations, and flexible adaptation to new relations without re-training).
3 Problem Description
Relation extraction: The task of relation extraction from the documents is typically defined as follows (Wang et al., 2019; Eberts and Ulges, 2021; Xu et al., 2021; Zhang et al., 2021; Zhou et al., 2021; Giorgi et al., 2022; Lu et al., 2022b; Tan et al., 2022; Wang Xu and Zhao, 2022; Xu et al., 2023). Given is a set of documents $\mathcal{D}=\{d_{i}\}_{i=1}^{M}$, where $M$ is the number of documents in the set. For each document $d_{i}$, the aim is to enumerate knowledge triplets $\{(r_{im},s_{im},o_{im})\}_{m=1}^{R_{i}}$, where $r_{im}\in\mathcal{R}$ is a relation and $s_{im}$ and $o_{im}$ are the subject and object of the relation $r_{im}$, and where $R_{i}$ is the number of relations in $d_{i}$. For instance, the document “The Reality Dysfunction is a science fiction novel by British writer Peter F. Hamilton normal-$\ldots$” yields the knowledge triplets (author, The Reality Dysfunction, Peter F. Hamilton), (genre, The Reality Dysfunction, science fiction), etc.
Difference to earlier works: Earlier works (see Sec. 2) generally address the above task through an intermediate step that involves named entities detection. Specifically, the aforementioned works first detect the named entities of a document $d_{i}$, i. e., $\{e_{ij}\}_{j=1}^{N_{i}}$, where $N_{i}$ is the number of entities in $d_{i}$. Then they proceed by predicting the relation(s) between each named entity pair $(e_{ij},e_{ij^{\prime}})_{j,j^{\prime}\in\{1,\ldots,N_{i}\},j\neq j^{\prime}}$ among the $R$ relations, where $e_{ij}$ is the subject and $e_{ij^{\prime}}$ is the object of predicted relation(s). As such, the number of predictions scales with the number of named entity pairs, i.e., it is in $O(N_{i}^{2})$. For the prediction, state-of-the-art methods make use of pre-trained LMs for which a multi-class classification head is fine-tuned. The classification head also predicts when there is no relation between the given pair and then discards the knowledge triplet. All remaining relations give the knowledge triplets $\{(r_{im},s_{im},o_{im})\}_{m=1}^{R_{i}}$.
In-context few-shot learning in REPLM: Our REPLM framework addresses the above drawbacks and approaches relation extraction as a triplet generation task. In this setup, the LM learns how to generate subject(s) and object(s) of a given relation from its in-context few-shot examples. Therefore, our REPLM framework does not require annotations of named entities. Our setup also facilitates the flexibility of adding new relations, simply by leveraging the given context examples. Specifically, for a given document $d_{i}\in\mathcal{D}$ and relation $r\in\mathcal{R}$, we prompt a pre-trained LM to generate the knowledge triplets of a relation $\{(r_{im},s_{im},o_{im})\,|\,r_{im}=r\}$ with no further fine-tuning.
In our REPLM framework, we have two sets of documents as input: (1) a distantly-supervised set $\mathcal{D}^{\text{dist}}$ for providing in-context few-shot examples and (2) a training set $\mathcal{D}^{\text{train}}$ for calibrating hyperparameters (which is optional). Details are in the next section.
4 Proposed REPLM Framework
Figure 1: Overview of our REPLM. Our framework takes a new document and relation as input and then proceeds along three steps: (1) selects a candidate pool of $N$ in-context examples; (2) constructs $L$ sets of such in-context examples; and (3) calculates the joint probabilities of subject-object pairs to extract knowledge triplets. Legend: subjects and objects are colored in blue and orange, respectively.
Approach (see Fig. 1): At a high-level, our framework seeks to infer the correct knowledge triplets $(r,s,o)$ from a given document $d_{i}$ and for a given relation $r$. To do so, we estimate the joint probability of a subject-object pair $(s,o)$ conditional on $d_{i}$ and $r$, i.e., $p(s,o\,|\,d_{i},r)$. After having estimated the probability, we simply rank the candidate subject-object pairs according to their probabilities and keep the top-ranked pairs as knowledge triplets. In our framework, we follow this approach but, as a main innovation, leverage a pre-trained LM to approximate $p(s,o\,|\,d_{i},r)$.
Learning via in-context few-shot examples: Pre-trained LMs are not explicitly trained for our relation extraction task, although they generally have the ability to extract information from a given context when guided properly. In our framework, we intentionally avoid the use of fine-tuning a pre-trained LM due to high computational cost and the inability of handling new relations. Instead, we provide guidance for our task via in-context few-shot examples. These examples demonstrate how to extract the subject-object pairs of relation $r$ from the given context. As a result, we can approximate $p(s,o\,|\,d_{i},r)\sim p(s,o\,|\,C,d_{i},r)$, where $C$ represents the selected set of in-context examples.
However, selecting only a single set of in-context examples may lead to a poor approximation of the probability $p(s,o\,|\,d_{i},r)$, because the selected in-context examples introduce bias in output generation (e. g., recency bias, label space of the in-context examples) as studied in prior literature (Liu et al., 2022a; Min et al., 2022b; Rubin et al., 2022; Hongjin et al., 2023; Wei et al., 2023). Instead, we mitigate the above bias by considering multiple sets of in-context examples. As a result, we calculate the joint probability of a subject-object pair as
$p(s,o\,|\,d_{i},r)=\sum_{l=1}^{L}p(C_{l}\,|\,d_{i},r)\cdot p(s,o\,|\,C_{l},d_{i},r),$ (1)
where we aggregate the outputs from $L$ sets of in-context examples. Here, $p(C_{l}\,|\,d_{i},r)$ is the weight of set $C_{l}$ of in-context examples, which measures how well $C_{l}$ is a candidate set compared to other sets of in-context examples.
Steps: Our REPLM framework proceeds along three steps: (1) it first selects a candidate pool for the in-context examples (Sec. 4.1); (2) it then constructs multiple sets of in-context examples and assigns their weights via a tailored approach (Sec. 4.2); and (3) it calculates the joint probabilities subject-object pairs to extract the knowledge triplets (Sec. 4.3). We describe the steps in the following.
4.1 Selecting Candidates for In-Context Examples
We now create a candidate pool of in-context few-shot examples for a given document $d_{i}$. Crucially, we generate our candidate pool in a way that, on the one hand, it is created via distant supervision and thus without human annotation, and, on the other hand, it is semantically related to the document $d_{i}$, thereby providing meaningful guidance.
Distant supervision: We create the in-context few-shot examples from the set $\mathcal{D}^{\text{dist}}$ generated by distant supervision. Specifically, $\mathcal{D}^{\text{dist}}$ is a dataset without any human annotation. In our implementation, we use the distantly-supervised split of DocRED (Yao et al., 2019), automatically created via an external knowledge base (KB). Reassuringly, we emphasize that this split comprises documents and knowledge triplets but it was created without any human annotation.
Distant supervision assumes that, if a document $\tilde{d}\in\mathcal{D}^{\text{dist}}$ contains both the subject and object of a knowledge triplet from a KB, it likely discusses their relationship. This premise allows for the automatic generation of annotated document sets. Such automated annotation introduces noisy labels (Yao et al., 2019; Cabot and Navigli, 2021), yet that can oftentimes be effective (Min et al., 2022b). A key benefit is that distantly supervised documents offer rich insights into label space, textual distributions, and expected output formats, over which the in-context few-shot learning paradigm in our REPLM can generalize. [^2] We further compare distant supervision with human-annotated data. They have the same performance, confirming the effectiveness of this approach for relation extraction (see Appendix D).
Semantic filtering: (i) We first filter the documents in $\mathcal{D}^{\text{dist}}$ so that we only keep the documents that contain at least one knowledge triplet of a relation $r$. We denote the result by $\mathcal{D}_{r}^{\text{dist}}$. Formally, $\mathcal{D}_{r}^{\text{dist}}$ is defined as
$\mathcal{D}_{r}^{\text{dist}}=\{d_{j}\,|\,\exists\,r^{\prime},s,o\ \mathrm{s.t.}\ (r^{\prime},s,o)\in d_{j}\\ \land r^{\prime}=r\land d_{j}\in\mathcal{D}^{\text{dist}}\}.$
(ii) We then retrieve $N$ documents from $\mathcal{D}_{r}^{\text{dist}}$ that are semantically most similar to $d_{i}$. For this, we leverage the technique from (Liu et al., 2022a) and encode the document $d_{i}$ and all the documents $\{d_{j}\,|\,d_{j}\in\mathcal{D}_{r}^{\text{dist}}\}$ into their embeddings via encoder $F_{\theta}$. (iii) We calculate the cosine similarity between the embeddings of $d_{i}$ and $d_{j}$, i.e., $\frac{F_{\theta}(d_{i})\cdot F_{\theta}(d_{j})}{||F_{\theta}(d_{i})||_{2}\cdot||F_{\theta}(d_{j})||_{2}}$. (iv) We keep the top-$N$ documents in $\mathcal{D}_{r}^{\text{dist}}$ in terms of cosine similarity to $d_{i}$ in embedding space. The selected $N$ documents form the context pool $\mathcal{D}_{r}^{\text{pool}}$, from which we construct multiple sets of in-context examples in the following.
4.2 Constructing Multiple Sets of In-Context Examples
For robustness, we create $L$ sets with in-context examples from our candidate pool $\mathcal{D}_{r}^{\text{pool}}$. We achieve this by random sampling of $K$ documents from $\mathcal{D}_{r}^{\text{pool}}$ across $L$ repetitions. As a result, we obtain $L$ sets of in-context examples, i.e., $C_{1},\ldots,C_{L}$. Then, we perform weighting at the set level.
Weighting at set level: The output from each context set $C_{l}$ should contribute to the relation extraction task proportional to some weight $p(C_{l}\,|\,d_{i},r)$. We calculate the weight as follows. First, we get a score for $C_{l}$ which is the average cosine similarity between the documents in $C_{l}$ and $d_{i}$, i. e.,
$\mathrm{score}(C_{l})=\frac{1}{K}\sum_{d_{j}\in C_{l}}\frac{F_{\theta}(d_{i})\cdot F_{\theta}(d_{j})}{||F_{\theta}(d_{i})||_{2}\cdot||F_{\theta}(d_{j})||_{2}}.$ (2)
We then use the score of $C_{l}$ to calculate the weight via
$p(C_{l}\,|\,d_{i},r)=\frac{\mathrm{exp}(\mathrm{score}(C_{l})/\tau)}{\sum_{l^{\prime}=1}^{L}\mathrm{exp}(\mathrm{score}(C_{l^{\prime}})/\tau)},$ (3)
where $\tau>0$ is for temperature scaling. Hence, $p(C_{l}\,|\,d_{i},r)$ represents how much the final output of REPLM should attend to the output generated from the context set $C_{l}$.
4.3 Computing Knowledge Triplet Probabilities
We now calculate the probabilities for subject-object pairs and then extract the knowledge triplets.
Prompting: We prompt our pre-trained LM with both (i) the in-context few-shot examples derived from $C_{l}$ and (ii) the document $d_{i}$ at the end of the prompt. For this, we first prepare the in-context demonstrations for each context set $C_{l}$. That is, we concatenate the documents $d_{j}$ in $C_{l}$, where each document $d_{j}$ is appended with its corresponding knowledge triplets $\{(r_{jm},s_{jm},o_{jm})\,|\,r_{jm}=r\}$. Each knowledge triplet is added in a new line (see Fig. 1). For the textual prompt, we separate the relation, subject, and object of $(r,s,o)$ with a special separator symbol <S>. This facilitates easier parsing of the subjects and objects generated.
Calculation of joint probability: We first obtain the log probabilities of both subject and object tokens under our pre-trained LM. We normalize the log probabilities by the length (i. e., number of tokens) of the subject and object. Formally, we compute (here: we directly write the exponent of the average log. probabilities for the ease of reading):
$p(s\,|\,C_{l},d_{i},r)$ =$\mathrm{len}(s)]{\prod_{k=1}^{\mathrm{len}(s)}p(s_{k}\,|\,s_{<k},C_{l},d_{i},r)},$ (4)
$p(o\,|\,s,C_{l},d_{i},r)$ =$\mathrm{len}(o)]{\prod_{k=1}^{\mathrm{len}(o)}p(o_{k}\,|\,o_{<k},s,C_{l},d_{i},r)},$ (5)
where $\mathrm{len}(s)$ and $\mathrm{len}(o)$ are the number of tokens of the subject and object, respectively. Afterward, we compute the joint probability $p(s,o\,|\,C_{l},d_{i},r)=p(s\,|\,C_{l},d_{i},r)\cdot p(o\,|\,s,C_{l},d_{i},r)$.
Ranking: As the final step, we calculate $p(s,o\,|\,d_{i},r)$ by aggregating over the context sets $C_{l}$, $l=1,\ldots,L$, as in Eq. (1) and repeat this for all generated subject-object pairs. We keep all generated knowledge triplets whose probability exceeds a certain threshold $\theta$, i. e., $\{(r,s,o)\,|\,p(s,o\,|\,d_{i},r)>\theta\}$. Of note, if a subject-object pair is not generated from a context set $C_{l}$, then $p(s,o\,|\,C_{l},d_{i},r)=0$.
Note that the latter step is different from state-of-the-art methods as these methods must enumerate over all possible subject-object pairs. Further, as can be seen here, our framework does not require named entities as input, which is another salient difference to many of the existing works.
5 Experimental Setup
We evaluate our framework using the DocRED (Yao et al., 2019), the largest document-level relation extraction dataset publicly available. DocRED includes 96 relation types and comprises three sets: (1) a distantly-supervised set with 101,873 documents, (2) a human-annotated train set with 3053 documents, and (3) a human-annotated development set with 998 documents. We provide further details about the DocRED dataset in the Appendix B. Importantly, in our experimental setup, we use distantly-supervised set for in-context few-shot learning ($\mathcal{D}^{\text{dist}}$) and evaluate the performance on the development set. Thereby, we ensure that our framework is solely trained without human annotation. To better understand the performance of our REPLM, we thus later also perform additional experiments (see Sec. 7) using sentence-level relation extraction datasets.
Baselines. We evaluate our framework against state-of-the-art methods for relation extraction that scale to document-level and, for comparability, do not require named entity recognition pipelines (see Table 1). These are: (1) REBEL (Cabot and Navigli, 2021) applies triplet linearization to extract the relations from the given document. (2) We also compare against REBEL-sent, which extracts the relations from a document in sentence-by-sentence manner. We include this variant because REBEL is originally trained at sentence level and, as we show later, it is the best-performing baseline method for the sentence-level relation extraction. Note that REBEL is the only baseline from the literature that does not require a named entity recognition pipeline for document-level relation extraction.
In our dataset, we use REBEL-large from Hugging Face [^3] https://huggingface.co/Babelscape/rebel-large, which is pre-trained by a tailored REBEL dataset [^4] https://huggingface.co/datasets/Babelscape/rebel-dataset. We note that REBEL-large is further fine-tuned on the human-annotated training set of DocRED, which may give it an (unfair) advantage. Hyperparameter selection and early-stopping are done based on some portion of the development set. This is again an (unfair) advantage for REBEL, whereas our REPLM does not need such dev set.
REPLM variants: We compare two variants of our proposed framework: (1) REPLM is the original variant as described above. Therein, we use fixed parameters. We also report a sensitivity analysis in Appendix J, where we vary the parameters to see that our performance is robust to different choices. (2) REPLM (params adj) is a variant for which the hyperparameters (e. g., temperature, threshold) are selected based on the training set.
Ablations: We perform ablation studies to assess the contribution of different components in our framework. Our ablations are: (1) REPLM (random fixed) randomly selects a single set of $K$ documents for each relation [^5] We also explored finding the “best” documents (Appendix K) of a relation. It requires evaluation against human-annotation and still performs worse than REPLM. Hence, we exclude its results.. However, the set is fixed across all evaluations. (2) REPLM (random all) randomly selects a set of $K$ documents for each relation and for each evaluation. (3) REPLM (best $\ominus$) selects the top-$K$ documents for each relation and each development document according to the cosine similarity. In-context examples are ordered from most similar to the least similar. (4) REPLM (best $\oplus$) similarly selects top $K$ documents for each relation and each development document. This time, in-context examples are ordered in reverse order, from least similar to most similar. We include these two alternatives to evaluate the effect (if any) of recency bias (Lu et al., 2022a; Hongjin et al., 2023).
Evaluation. We calculate the F1 score for each relation, counting an extraction as correct only if the subject and object exactly align with the ground-truth. Thus, extracted relations missing in the development set are false positives, while those in the set but not generated are false negatives.
Implementation. We use GPT-JT [^6] https://huggingface.co/togethercomputer/GPT-JT-6B-v1 ($\sim$6B parameters) as our pre-trained LM for in-context few-shot learning. Appendix E provides all details of our REPLM framework.
6 Results
6.1 Overall Performance
Figure 2: F1 scores per relation type (darker = better). Missing color means that no correct predictions were made for this relation. F1 scores are normalized by the maximum value for each relation. Relations are in decreasing order of their number of knowledge triplets.
First, we evaluate how accurately our REPLM framework extracts the relations from the given documents by comparing them against human annotations (Fig. 2) [^7] F1 scores on each relation are given in Appendix H.. Overall, our REPLM and REPLM (params adj) achieve state-of-the-art performance on most relation types. This pattern is especially pronounced for relations with a large number of knowledge triplets (e.g., P17: country, P131: located in, P27: country of citizenship).
Table 2: Document-level relation extraction results. Shown: Micro F1 scores.
Method F1 score
REBEL (Cabot et al., 2021) 26.17
REBEL-sent (Cabot et al., 2021) 27.52
REPLM (random fixed) 21.04  $\pm$  0.17
REPLM (random all) 21.14  $\pm$  0.09
REPLM (best $\ominus$) 31.31
REPLM (best $\oplus$) 31.04
REPLM (ours) 33.93
REPLM (params adj) (ours) 35.09
Higher is better. Best value in bold.
Table 2 reports the overall performance, i.e., the micro F1 score over all relation types. Our REPLM achieves an F1 score of 33.93, and our REPLM (params adj) an F1 score of 35.09. The slight advantage of the latter is expected and can be attributed to the additional hyperparameter tuning. For comparison, the REBEL-sent baseline registers only an F1 score of 27.52. In sum, our framework performs the best and results in an improvement of +27 %. Note that REBEL was even fine-tuned on some samples of the dev set, which again demonstrates the clear superiority of our framework. We observe that REPLM outputs, on average, 20.21 knowledge triplets per document while REBEL outputs only 4.93; we discuss the implications later.
In our ablation study, we compare different variants of our REPLM to understand the source of performance gains (see Fig. 2 and Table 2). We make the following observations: (1) Retrieving the best in-context examples improves the performance compared to random examples by more than 48 % (REPLM (best $\ominus$) and REPLM (best $\oplus$) vs. REPLM (random fixed) and REPLM (random all)). (2) We do not observe that a recency bias plays a decisive role for our results, as both REPLM (best $\ominus$) and REPLM (best $\oplus$) reach a similar performance. (3) Our complete framework brings a significant improvement over REPLM (best $\ominus$) and REPLM (best $\oplus$) (+18 %) by aggregating multiple sets of most relevant in-context examples, thus establishing the importance of using multiple sets.
Insights. We conjecture that our REPLM extracts more relations than REBEL, as it further identifies the missing annotations in DocRED. In Appendix F, we empirically validate that, for each relation type, there are dev documents with no annotation, which are semantically similar to the documents containing at least one knowledge triplet of a relation. It suggests that these dev documents potentially include the relation but lack the annotation. To confirm this conjecture, we perform a manual validation that we examine cases where our REPLM fails. We find that many relations extracted by our method are actually correct but considered as false positives due to missing human annotations. For example, our REPLM generates the relation (author, Chaosmosis, Félix Guattari) but it is not annotated and therefore marked as incorrect (see right part of Fig. 1). We provide additional examples in the Appendix G.
6.2 Comparison Against External Knowledge
Figure 3: F1 scores per relation type (darker = better). Here, the comparison is made against annotations that additionally make use of external knowledge and should thus more closely reflect the ground-truth. Relation types are arranged in decreasing order of their number of knowledge triplets. For visibility, F1 scores are normalized by the maximum value for each relation.
The above evaluations were constrained by relying solely on the human annotations on DocRED, potentially penalizing accurate methods due to missing annotations. We now repeat our evaluations using an alternative gold standard for a more comprehensive benchmark.
Table 3: Document-level relation extraction results evaluated via external KB. Shown: Micro F1 scores.
Method F1-Score
REBEL (Cabot et al., 2021) 20.30
REBEL-sent (Cabot et al., 2021) 20.00
REPLM (ours) 32.33
REPLM (params adj) (ours) 36.51
Higher is better. Best value in bold.
Ground-truth via external knowledge: To locate the missing annotations in DocRED, we aggregate all relations extracted from all methods on all documents. We then check the correctness of the extracted relations via an external KB. Specifically, we leverage the pipeline from HELM (Liang et al., 2022) and check if the generated knowledge triplet exists in Wikidata (Vrandečić and Krötzsch, 2014). We add all the matched triplets to the existing list of ground-truth triplets from DocRED and repeat the evaluation. As a result, total number of relations in the development set increased from 12,212 to 18,592.888We note that the increase in the number of relations does not necessarily imply an improvement in the F1 score for our REPLM. The extracted relations are still filtered by the probability threshold $\theta$, which, in turn, reduces the recall (and possibly the F1 score).
Results: For our new dataset with external ground-truth, our framework outperforms REBEL by a considerable margin across most relation types (Fig. 3) [^9] The details of evaluation via external KB are in Appendix I.. This is also seen in the overall performance (Table 3). For example, our REPLM improves the F1 score over REBEL by more than 59 % (32.33 vs. 20.30). The improvement for REPLM (params adj) is even larger and amounts to 80 % (36.51 vs 20.30).
7 Ablation Study: Performance for Sentence-Level Relation Extraction
After showing the effectiveness of our REPLM in document-level relation extraction, we additionally perform sentence-level relation extraction as an ablation study, albeit not our primary benchmarking focus. Despite this, our REPLM still matches or surpasses the performance of fine-tuned LMs.
Datasets. We conduct extensive experiments on three sentence-level relation extraction benchmark datasets, namely CONLL04 (Roth and Yih, 2004), NYT (Riedel et al., 2010), and ADE (Gurulingappa et al., 2012). We provide the details of each dataset in Appendix C.
Baselines. We compare REPLM against fine-tuned LMs that achieve state-of-the-art results: SpERT (Eberts and Ulges, 2020), Table-sequence (Wang and Lu, 2020), TANL and TANL (multi-dataset) [^10] This version is jointly trained on all 3 datasets. (Paolini et al., 2021), REBEL (w/o pre-training) and REBEL (Cabot and Navigli, 2021) [^11] The same REBEL from Sec. 6 that is pre-trained by an external dataset before fine-tuning.
Performance. Table 4 shows the micro F1 scores of each method on CONLL04, NYT, and ADE datasets. Our REPLM achieves the best performance on ADE against all the supervised methods with an F1 score of 82.5. We observe that our REPLM performs on par with the state-of-the-art methods on CONLL04 dataset with an F1 score of 72.9, while REBEL attains the best score via an external dataset. For NYT, our REPLM achieves a lower score than state-of-the-art methods. The reason is, the entire NYT is curated via distant supervision, which results in noisy relations. As the baseline methods are fine-tuned with the same dataset, they can memorize the same knowledge triplets from the training set, and therefore, achieve spuriously high performance. The detailed explanation is in Appendix C.
The above analysis also justifies our choice of using REBEL, which is the best of the baselines. It is further the only model that can extract relations from the documents without named entity annotation. Hence, the above ablation also explains why we choose REBEL as the main baseline in Sec. 6.
Table 4: Sentence-level relation extraction. Shown: Micro F1.
Method CONLL04 NYT ADE
SpERT (Eberts and Ulges, 2020) 71.5 _ 79.2
Table-sequence (Wang and Lu, 2020) 73.6 _ 80.1
TANL (Paolini et al., 2021) 71.4 90.8 80.6
TANL (multi-dataset) (Paolini et al., 2021) 72.6 90.5 80.0
REBEL (w/o pre-training) (Cabot et al., 2021) 71.2 91.8 81.7
REBEL (Cabot et al., 2021) 75.4 92.0 82.2
REPLM (ours) 72.9 81.0 82.5
Higher is better. Best value in bold.
Further insights. What is the effect of the number of in-context examples? We repeat the same experiment on CONLL04 with a varying number of in-context examples. Fig. 3(a) shows the F1 score on each relation and the overall score by varying $K$ from 3 to 11. We observe that, in general, more in-context examples yield better F1 scores. Informed by this observation, we used the highest number of in-context examples that fit into the context window for our experiments, which is $K=5$ for document-level relation extraction. Detailed results are given in Appendix L.
(a) Performance with varying
num. of in-context examples.
(b) Change in performance with
random entity names.
Figure 4: Ablation studies on CONLL04.
Is REPLM actually learning to extract relations? Or does it only retrieve facts from memory? We design a novel experiment to identify whether our REPLM is learning to extract relations from the input text or it is simply retrieving the facts from its memory. To the best of our knowledge, we are the first to shed more light on the models’ learning ability for the relation extraction task. For this experiment, we replaced all the entities with random names in CONLL04 dataset (for both training and test set) that are not mentioned anywhere in the web. Fig. 3(b) compares the performance against the original dataset. The overall performance decreases only slightly when using the random entities (F1 score of 70.47 vs. 72.9), which is still on par with the state-of-the-art. Therefore, it confirms that our REPLM is an effective method in learning to extract the relations from the context. We provide the experiment details and elaborate on the reasons of slight performance decrease in Appendix M.
8 Discussion
Benefits: Our REPLM framework offers many benefits in practice: (1) REPLM eliminates the need for named entity recognition pipelines in our task and thus the noise that comes with it; (2) REPLM does not require human annotations but leverages in-context few-shot learning; and (3) REPLM offers great flexibility as it allows to incorporate new relations without re-training.
Implications: Our findings have direct implications for constructing datasets and benchmarking methods aimed at relation extraction from the textual documents. Evidently, earlier datasets such as DocRED (Yao et al., 2019), the largest dataset for document-level relation extraction publicly available, lack comprehensiveness and thus miss important – but correct – annotations. This may penalize correct methods and instead favor overly simple output during benchmarking. Hence, it will be important for future research to develop new datasets and effective evaluation paradigms.
Broader Impact Statement
This paper presents work whose primary goal is to advance the field of machine learning. We would like to emphasize the positive and negative implications that underlie machine learning more broadly. Our REPLM may help to fill the gaps in existing knowledge bases. It is likely that such gaps are particularly dominant for marginalized groups, so that our work can have a positive impact for improving coverage in knowledge bases for diverse groups. Still, LMs are far from perfect as their performance may vary. Hence, careful and responsible use is indicated, especially when filling in knowledge bases with important societal or ethical implications or sensitive content more broadly.
Appendix A Related Work
In-context few-shot learning of LMs: In-context few-shot learning has been widely adopted in text classification tasks (Holtzman et al., 2021; Liu et al., 2022a; Lu et al., 2022a; Min et al., 2022a, b; Zhao et al., 2021) and further extended to other tasks such as question answering (Holtzman et al., 2021; Liu et al., 2022a; Min et al., 2022b), fact retrieval (Zhao et al., 2021), table-to-text generation (Liu et al., 2022a), and mapping utterances to meaning representations (Rubin et al., 2022). However, we are not aware of any earlier work that leveraged the in-context few-shot learning paradigm for document-level relation extraction.
LMs as knowledge bases: Research has focused on probing the knowledge in LMs. For example, Petroni et al. (2019) introduced the LAMA dataset, a dataset with cloze-style templates for different relations, which allows to probe factual knowledge in LMs. Many works have been introduced to achieve state-of-the-art results via prompt-tuning (Hao et al., 2022; Lester et al., 2021; Li and Liang, 2021; Liu et al., 2021b, 2022b; Newman et al., 2022; Perez et al., 2021; Poerner et al., 2020; Shin et al., 2020; Zhong et al., 2021). Yet, some works further find that, when evaluated as knowledge bases, LMs suffer from inconsistency (AlKhamissi et al., 2022; Elazar et al., 2021), learn shallow heuristics rather than facts (Elazar et al., 2022), have inferior performance in the long tail (Kandpal et al., 2022), and exhibit prompt bias (Cao et al., 2021).
However, we note that the above research stream is different from our work in two salient ways. (1) In the above research stream, LMs are prompted to retrieve factual knowledge from its memory, whereas we aim to extract the relational knowledge from the context. (2) In the above research stream, LM prompts are structured as “fill-in-the-blank” cloze statement. For example, the task is to output only the correct object, but where both the subject and relation are given. Instead of predicting only the object, our goal is to output the entire knowledge triplet. That is, subject, relation, and object must be inferred together.
Relation extraction via pattern-based and statistical methods: A detailed review of the different methods is provided in Pawar et al. (2017) and Weikum et al. (2021), while we only present a brief summary here. Early works extracted relations from text via pattern-based extraction methods. Specifically, these works introduced automated methods to extract textual patterns corresponding to each relation and each entity type (Carlson et al., 2010; Jiang et al., 2017; Nakashole et al., 2012). Their main limitation is that the automatically constructed patterns involve many mistakes, which, in turn, require human experts to examine and correct them (Han et al., 2020).
Another research stream focused on relation extraction via statistical methods. Examples are crafting custom features for relation classification (Jiang and Zhai, 2007; Nguyen et al., 2007), designing customized kernels for support vector machines(Nguyen et al., 2007; Wang, 2008; Zhang et al., 2006b, a), graphical modeling and inference of relations (Sarawagi and Cohen, 2004; Yu and Lam, 2010), and leveraging knowledge graph embeddings for relation prediction (Lin et al., 2015; Wang et al., 2014).
However, the above methods have only a limited capacity in capturing complex interactions between entities, as compared to state-of-the-art neural networks (Han et al., 2020). On top of that, both pattern-based and statistical methods require large datasets with human annotation for training.
Relation extraction via neural networks: Initial methods for relation extraction based on neural network approaches made use of convolutional neural network (CNN) (Zeng et al., 2014) and long short-term memory (LSTM) (Zhou et al., 2016) architectures. These works process the pre-computed word embeddings and then classify the relation for the given named entity pair. Follow-up works proposed joint learning of entity extraction and relation classification, again via CNN (Adel and Schütze, 2017; Zheng et al., 2017) and LSTM (Katiyar and Cardie, 2017; Miwa and Bansal, 2016) architectures. However, these models are not flexible enough to model the complex interactions between named entities to classify the relation, as compared to LMs.
Relation extraction via LMs: State-of-the-art methods for relation extraction are based on fine-tuning pre-trained LMs. Specifically, these methods use pre-trained LMs such as BERT (Devlin et al., 2019), RoBERTa (Liu et al., 2019), and SciBERT (Beltagy et al., 2019) and fine-tuned them for relation extraction. For instance, Wang et al. (2019) fine-tuned BERT to classify the relation between each named entity pair in a given sentence. There have been various follow-up works to improve performance by learning complex dependency between named entities. To achieve this, Wang and Lu (2020) jointly trained LSTM and BERT to get two distinct representations of the entities; Zhang et al. (2021) further incorporate semantic segmentation module into the fine-tuning of BERT; Zhou et al. (2021) propose adaptive thresholding and localized context pooling; Xu et al. (2021) explicitly model the dependencies between entity mentions; Paolini et al. (2021) augmented the original sentences with the entity and relation types; Tan et al. (2022) use axial attention module for learning the interdependency among named entity pairs; Wang Xu and Zhao (2022) propose sentence importance estimation; Xiao et al. (2022) include additional tasks such as coreference resolution, entity typing, and evidence retrieval; Xu et al. (2023) improves the model performance via synthetic data generation; Hu et al. (2023) incorporates rationale extraction from the sentence; and Zhang et al. (2023a) leverages self-distillation to facilitate relational reasoning. However, all of these works require the named entities to be annotated and provided as input at both training and test time.
Appendix B DocRED Dataset
For our experiments, we use DocRED (Yao et al., 2019), the largest publicly available dataset for document-level relation extraction. We provide the detailed statistics of each relation type in Tables 5 and 6 (note: the different columns compare the different subsets for distant supervision, human-annotated training, and human-annotated dev).
Pre-processing. The original documents in the DocRED dataset are provided only in a tokenized format, e. g., the document is represented as a list of token, where each punctuation mark and word is a different token. We follow the earlier works (Cabot and Navigli, 2021; Yao et al., 2019) and concatenate the tokens with a white space in between to construct the entire document. This approach may introduce typos in the documents; for instance, the original text “Tarzan’s Hidden Jungle is a 1955 black-and-white film ...” is reconstructed as “Tarzan ’s Hidden Jungle is a 1955 black - and - white film ...”. We initially tried to fix these typos via spelling correction libraries, such as FastPunct [^12] https://pypi.org/project/fastpunct/, but later found that the typos are propagated to the labels, which may impede performance and eventually comparability of our results. Therefore, we decided to follow the same pre-processing as earlier works, as it allows us to operate on the same labels as in earlier work and thus ensures comparability of our results. Sec. G shows some examples of documents after the pre-processing step.
Table 5: DocRED statistics
Relation ID Relation Name # Docs in Dist. Sup. # Relations in Dist. Sup. # Docs in Train # Relations in Train # Docs in Dev # Relations in Dev
P6 head of government 4948 6859 133 210 38 47
P17 country 68402 313961 1831 8921 585 2817
P19 place of birth 21246 31232 453 511 135 146
P20 place of death 15046 24937 170 203 50 52
P22 father 5287 9065 164 273 41 57
P25 mother 1828 2826 50 74 10 15
P26 spouse 4327 9723 134 303 34 74
P27 country of citizenship 45553 126360 1141 2689 384 808
P30 continent 7247 18792 121 356 38 121
P31 instance of 3790 5561 74 103 34 48
P35 head of state 3127 4257 87 140 32 51
P36 capital 27621 34047 66 85 24 27
P37 official language 4040 6562 82 119 29 47
P39 position held 982 1692 15 23 6 8
P40 child 5794 11831 177 360 45 81
P50 author 5265 8856 162 320 49 93
P54 member of sports team 2693 12312 80 379 36 166
P57 director 5891 9865 153 246 58 90
P58 screenwriter 4680 7952 83 156 24 35
P69 educated at 5201 8413 220 316 63 92
P86 composer 2778 4249 44 79 21 57
P102 member of political party 5464 11582 191 406 51 98
P108 employer 4168 6775 126 196 30 54
P112 founded by 5856 7700 74 100 20 27
P118 league 2142 6024 63 185 29 56
P123 publisher 2426 $4$ 81 172 29 69
P127 owned by 4907 7554 91 208 36 76
P131 located in the administrative territorial entity 44307 143006 1224 4193 389 1227
P136 genre 982 1948 34 111 7 14
P137 operator 1982 3011 52 95 18 41
P140 religion 2515 5143 60 144 26 82
P150 contains administrative territorial entity 34615 62646 1002 2004 310 603
P155 follows 8360 12236 117 188 43 69
P156 followed by 7958 11576 120 192 38 51
P159 headquarters location 12653 17089 206 264 57 86
P161 cast member 6575 21139 163 621 62 226
P162 producer 4434 6739 77 119 32 50
P166 award received 2852 6322 105 173 35 66
P170 creator 3485 6036 96 231 25 40
P171 parent taxon 860 2167 28 75 6 17
P172 ethnic group 6022 7563 63 79 24 30
P175 performer 10783 27945 344 1052 101 332
P176 manufacturer 1260 2737 27 83 9 40
P178 developer 2403 6368 73 238 30 75
P179 series 2404 3800 72 144 27 63
P190 sister city 3388 11471 2 4 1 2
P194 legislative body 2863 2989 136 166 36 56
P205 basin country 2249 3299 61 85 21 32
Table 6: DocRED statistics (continued)
Relation ID Relation Name # Docs in Dist. Sup. # Relations in Dist. Sup. # Docs in Train # Relations in Train # Docs in Dev # Relations in Dev
P206 located in or next to body of water 3859 6585 109 194 35 83
P241 military branch 1589 2633 69 108 30 42
P264 record label 4524 14804 154 583 49 237
P272 production company 1417 2151 49 82 19 36
P276 location 5281 6654 130 172 55 74
P279 subclass of 1822 2736 39 77 19 36
P355 subsidiary 1761 2436 51 92 18 30
P361 part of 17335 28245 382 596 119 194
P364 original language of work 1061 2274 32 66 11 30
P400 platform 1565 5825 52 304 14 69
P403 mouth of the watercourse 1700 2475 49 95 19 38
P449 original network 2953 4237 97 152 20 39
P463 member of 7364 15272 208 414 55 113
P488 chairperson 1792 2216 49 63 15 21
P495 country of origin 17160 36029 300 539 112 212
P527 has part 13318 22596 317 632 94 177
P551 residence 2629 3197 25 35 5 6
P569 date of birth 26474 33998 893 1044 286 343
P570 date of death 20905 28314 587 805 180 255
P571 inception 19579 26699 393 475 127 154
P576 dissolved, abolished or demolished 5064 7057 52 79 25 39
P577 publication date 17636 37538 576 1142 193 406
P580 start time 5374 6549 96 110 30 32
P582 end time 4943 6144 47 51 18 23
P585 point in time 2457 2920 80 96 29 39
P607 conflict 4119 8056 114 275 46 114
P674 characters 1594 3447 62 163 25 74
P676 lyrics by 1677 2415 30 36 5 8
P706 located on terrain feature 3157 5063 74 137 29 60
P710 participant 2839 4985 95 191 22 57
P737 influenced by 1166 2071 9 9 3 10
P740 location of formation 3885 4531 53 62 12 15
P749 parent organization 2425 3335 47 92 27 40
P800 notable work 4053 5275 102 150 32 56
P807 separated from 1438 2210 2 2 1 2
P840 narrative location 2026 2573 38 48 11 15
P937 work location 5063 7470 69 104 19 22
P1001 applies to jurisdiction 7471 9945 204 298 55 83
P1056 product or material produced 460 624 27 36 6 9
P1198 unemployment rate 1330 1622 2 2 1 1
P1336 territory claimed by 880 1600 18 33 6 10
P1344 participant of 1707 3574 87 223 28 57
P1365 replaces 1490 1811 13 18 9 10
P1366 replaced by 2214 2771 25 36 10 10
P1376 capital of 25241 29816 62 76 20 21
P1412 languages spoken, written or signed 2781 6313 91 155 24 46
P1441 present in work 2872 6763 88 299 34 116
P3373 sibling 3335 11123 102 335 26 134
Appendix C Details on Sentence-Level Relation Extraction
We select three sentence-level relation extraction datasets to show the effectiveness of our REPLM framework against the state-of-the-art supervised methods.
CONLL04 (Roth and Yih, 2004) is consisting of sentences collected from the news articles. The authors manually annotated the entities and 5 relation types for each sentence. Following the earlier literature, we used the same splits as Eberts and Ulges (2020). The detailed statistics for each relation type are given in Table 7.
Table 7: CONLL04 statistics
Relation Name # Sentences in Train # Relations in Train # Sentences in Validation # Relations in Validation # Sentences in Test # Relations in Test
Kill 160 179 39 42 46 47
Live_In 270 326 68 84 82 98
Located_In 187 245 52 65 58 90
OrgBased_In 213 260 47 71 70 96
Work_For 208 250 57 69 65 76
Overall 922 1260 231 331 288 407
NYT (Riedel et al., 2010) is composed of sentences from New York Times, containing 24 relation types. The detailed statistics are given in Table 8.
It is important to note that the relations in this dataset are annotated via “distant supervision”, using the knowledge triplets from FreeBase (Bollacker et al., 2008). As a result, the evaluation on the test set becomes noisy. For instance, the sentence in the test set “Mr. Abbas, speaking before a meeting in Paris with the French president, Jacques Chirac, said he was sorry for the shootings on Sunday.” is annotated with the following knowledge triplet “(place of birth, Jacques Chirac, Paris)”, although the birthplace of Jacques Chirac cannot be inferred from the sentence.
We further found that the overlap of relations between train and test set is high. For the relation type place of birth, 166 out of 260 relations (i. e., the exact (relation, subject, object) triplet) in test set appear in the training set. Therefore, although the evaluation on the test is noisy, the baseline methods leverage the supervised training and they can memorize the relations from the train set at the test time. We hypothesize this as the main reason of the inferior performance of our REPLM framework on this dataset specifically, while achieving the state-of-the-art performance at all other evaluations.
Table 8: NYT statistics
Relation Name # Sentences in Train # Relations in Train # Sentences in Validation # Relations in Validation # Sentences in Test # Relations in Test
advisors 37 37 5 5 3 3
capital 6042 6121 557 567 649 659
child 407 437 45 46 32 40
contains_administrative_territorial_entity 4889 5111 462 497 496 527
country 4889 5111 462 497 496 527
country_of_citizenship 6136 6606 545 596 518 548
country_of_origin 19 19 1 1 1 1
denonym 29 32 3 3 1 1
employer 4546 4734 428 448 401 417
ethnicity 19 19 1 1 1 1
founded_by 649 682 53 58 58 61
headquarters_location 180 186 21 22 17 17
industry 1 1 _ _ _ _
location 37626 42961 3302 3818 3296 3835
location_of_formation 344 346 35 35 35 35
major_shareholder 229 238 21 21 31 32
member_of_sports_team 180 186 21 22 17 17
neighborhood_of 4329 4682 403 444 338 374
occupation 2 2 _ _ _ _
place_of_birth 2649 2703 215 217 256 260
place_of_death 1652 1676 125 128 127 131
religion 54 56 7 7 5 5
residence 5883 6182 506 531 570 597
shareholders 229 238 21 21 31 32
Overall 56196 88366 5000 7985 5000 8120
ADE (Gurulingappa et al., 2012) contains the sentences from biomedical domain and it has only one relation type, which is adverse effect. The original dataset contains 10 folds of train and test splits. Following the earlier work (Cabot and Navigli, 2021), we use the test set of the first fold for the evaluation. The statistics are given in Table 9.
Table 9: ADE statistics
Relation Name # Sentences in Train # Relations in Train # Sentences in Validation # Relations in Validation # Sentences in Test # Relations in Test
Adverse-Effect 3845 5980 _ _ 427 653
Appendix D In-Context Few-Shot learning based on Distant Supervision vs. Human Annotation
We perform an ablation study to observe the effect of using distantly-supervised vs. human-annotated documents as in-context few-shot examples. We perform such comparison on our four framework variants, namely REPLM (random fixed), REPLM (random all), REPLM (best $\ominus$), and REPLM (best $\oplus$). For methods with random in-context examples, the performance may be subject to variability across which seed is picked (whereas the performance is deterministic for the other methods), and, hence, we report the standard deviation for this subset of the methods by averaging the performance across 10 runs. Of note, to directly compare the impact of in-context examples, we deliberately considered our REPLM variants without aggregation over the multiples sets of in-context examples.
Tables 10 to 21 show the comparison between distant supervision vs. human annotation. For random in-context examples (i. e., REPLM (random fixed) and REPLM (random all)), distant supervision and human annotation performs at the same level. For retrieving the semantically most similar context examples (i. e., REPLM (best $\ominus$) and REPLM (best $\oplus$)), we observe cases where distant supervision actually improves the result (e. g., P6, P155, P179). However, the overall performance is largely similar. This confirms our choice of using distantly-supervised documents as in-context examples and eliminates the need for human annotation.
Table 10: Ablation study. Comparing the performance across the in-context examples from distant supervision vs. human-annotated train set. Shown are F1 scores on each relation. (Part 1 / 12)
Method Context Source P6 P17 P19 P20 P22 P25 P26 P27
REPLM (random fixed) Train 20.29  $\pm$  8.01 11.18  $\pm$  2.44 75.68  $\pm$  6.45 71.92  $\pm$  4.61 13.17  $\pm$  5.66 0.00  $\pm$  0.00 24.09  $\pm$  5.00 23.03  $\pm$  3.25
REPLM (random fixed) Dist. Sup. 24.28  $\pm$  10.53 11.34  $\pm$  4.48 66.67  $\pm$  12.71 49.80  $\pm$  16.07 9.76  $\pm$  4.90 0.00  $\pm$  0.00 26.73  $\pm$  4.38 22.41  $\pm$  2.53
REPLM (random all) Train 19.76  $\pm$  2.59 11.03  $\pm$  0.88 77.17  $\pm$  1.50 68.24  $\pm$  3.54 11.94  $\pm$  3.03 0.00  $\pm$  0.00 24.03  $\pm$  2.73 21.94  $\pm$  1.50
REPLM (random all) Dist. Sup. 25.90  $\pm$  2.44 11.05  $\pm$  0.46 68.49  $\pm$  3.09 48.35  $\pm$  4.82 9.51  $\pm$  2.86 0.00  $\pm$  0.00 26.74  $\pm$  3.83 23.26  $\pm$  1.39
REPLM (best $\ominus$) Train 18.18 19.62 79.29 69.90 15.09 7.14 33.33 28.03
REPLM (best $\ominus$) Dist. Sup. 35.96 24.60 71.38 50.39 18.18 15.38 33.33 29.14
REPLM (best $\oplus$) Train 18.18 19.65 74.20 72.00 20.00 5.71 30.43 28.27
REPLM (best $\oplus$) Dist. Sup. 35.96 24.02 68.63 62.50 17.70 7.41 22.97 28.33
Table 11: Ablation study. Comparing the performance across the in-context examples from distant supervision vs. human-annotated train set. Shown are F1 scores on each relation. (Part 2 / 12)
Method Context Source P30 P31 P35 P36 P37 P39 P40 P50
REPLM (random fixed) Train 16.94  $\pm$  6.32 0.00  $\pm$  0.00 26.86  $\pm$  3.49 16.56  $\pm$  4.57 29.69  $\pm$  10.49 0.00  $\pm$  0.00 15.13  $\pm$  6.04 28.38  $\pm$  4.20
REPLM (random fixed) Dist. Sup. 13.58  $\pm$  1.66 5.21  $\pm$  2.48 28.79  $\pm$  4.96 18.67  $\pm$  5.96 22.59  $\pm$  8.79 0.00  $\pm$  0.00 12.97  $\pm$  3.62 27.66  $\pm$  4.05
REPLM (random all) Train 15.43  $\pm$  2.93 6.66  $\pm$  1.82 27.40  $\pm$  3.48 19.92  $\pm$  8.34 27.73  $\pm$  3.60 0.00  $\pm$  0.00 21.05  $\pm$  3.09 28.01  $\pm$  3.89
REPLM (random all) Dist. Sup. 14.75  $\pm$  2.18 6.33  $\pm$  2.15 26.96  $\pm$  3.17 14.85  $\pm$  5.65 24.13  $\pm$  5.48 0.00  $\pm$  0.00 11.63  $\pm$  3.12 26.74  $\pm$  3.45
REPLM (best $\ominus$) Train 19.14 11.90 27.27 22.22 36.36 15.38 29.14 34.18
REPLM (best $\ominus$) Dist. Sup. 31.25 6.59 31.46 47.06 31.17 25.00 20.38 38.60
REPLM (best $\oplus$) Train 27.84 9.09 21.78 17.54 34.21 0.00 30.87 40.26
REPLM (best $\oplus$) Dist. Sup. 30.85 6.82 37.21 45.28 22.78 25.00 26.42 37.66
Table 12: Ablation study. Comparing the performance across the in-context examples from distant supervision vs. human-annotated train set. Shown are F1 scores on each relation. (Part 3 / 12)
Method Context Source P54 P57 P58 P69 P86 P102 P108 P112
REPLM (random fixed) Train 36.25  $\pm$  12.29 30.89  $\pm$  1.64 25.41  $\pm$  5.63 60.96  $\pm$  4.38 17.27  $\pm$  5.49 37.26  $\pm$  5.68 34.50  $\pm$  3.90 24.48  $\pm$  9.40
REPLM (random fixed) Dist. Sup. 42.71  $\pm$  10.02 34.55  $\pm$  4.59 27.65  $\pm$  6.61 53.00  $\pm$  9.62 16.16  $\pm$  8.57 30.82  $\pm$  9.53 33.29  $\pm$  4.49 14.30  $\pm$  8.81
REPLM (random all) Train 39.88  $\pm$  6.70 31.76  $\pm$  1.75 27.32  $\pm$  6.41 61.91  $\pm$  2.44 17.29  $\pm$  5.43 32.78  $\pm$  2.11 32.23  $\pm$  2.59 22.26  $\pm$  6.95
REPLM (random all) Dist. Sup. 43.80  $\pm$  3.48 32.63  $\pm$  3.59 27.32  $\pm$  3.10 57.18  $\pm$  3.74 13.23  $\pm$  4.60 32.85  $\pm$  3.32 32.67  $\pm$  5.06 16.06  $\pm$  6.68
REPLM (best $\ominus$) Train 48.67 33.70 35.48 67.07 30.59 37.66 41.38 12.77
REPLM (best $\ominus$) Dist. Sup. 48.30 47.62 34.38 57.47 23.26 44.44 30.19 32.65
REPLM (best $\oplus$) Train 50.57 40.72 29.41 59.63 35.16 39.74 35.96 17.02
REPLM (best $\oplus$) Dist. Sup. 40.93 43.43 39.34 58.29 40.45 40.48 34.29 34.78
Table 13: Ablation study. Comparing the performance across the in-context examples from distant supervision vs. human-annotated train set. Shown are F1 scores on each relation. (Part 4 / 12)
Method Context Source P118 P123 P127 P131 P136 P137 P140 P150
REPLM (random fixed) Train 33.05  $\pm$  6.31 24.29  $\pm$  3.37 10.23  $\pm$  3.89 15.66  $\pm$  2.99 27.79  $\pm$  8.24 14.49  $\pm$  4.58 8.38  $\pm$  3.52 21.52  $\pm$  3.58
REPLM (random fixed) Dist. Sup. 32.16  $\pm$  5.80 19.81  $\pm$  4.20 11.10  $\pm$  4.41 14.32  $\pm$  2.81 19.70  $\pm$  6.09 9.99  $\pm$  3.84 0.00  $\pm$  0.00 23.36  $\pm$  2.23
REPLM (random all) Train 32.13  $\pm$  5.64 26.43  $\pm$  3.86 13.17  $\pm$  2.31 15.58  $\pm$  0.62 25.20  $\pm$  7.89 11.77  $\pm$  3.69 6.73  $\pm$  2.76 21.93  $\pm$  1.81
REPLM (random all) Dist. Sup. 33.51  $\pm$  5.47 19.87  $\pm$  3.15 10.50  $\pm$  3.94 15.54  $\pm$  0.65 21.93  $\pm$  6.65 9.83  $\pm$  2.77 0.00  $\pm$  0.00 21.91  $\pm$  1.52
REPLM (best $\ominus$) Train 33.33 20.75 13.79 22.45 22.22 21.18 11.68 0.00
REPLM (best $\ominus$) Dist. Sup. 44.04 30.19 18.49 25.50 20.00 12.90 14.17 31.28
REPLM (best $\oplus$) Train 36.36 29.41 12.90 22.95 32.00 15.58 9.66 0.00
REPLM (best $\oplus$) Dist. Sup. 37.84 19.82 18.03 26.59 26.09 9.84 13.53 30.25
Table 14: Ablation study. Comparing the performance across the in-context examples from distant supervision vs. human-annotated train set. Shown are F1 scores on each relation. (Part 5 / 12)
Method Context Source P155 P156 P159 P161 P162 P166 P170 P171
REPLM (random fixed) Train 5.28  $\pm$  3.31 11.22  $\pm$  5.91 24.25  $\pm$  8.04 27.55  $\pm$  7.43 18.40  $\pm$  7.91 27.09  $\pm$  2.02 5.44  $\pm$  2.44 14.23  $\pm$  3.10
REPLM (random fixed) Dist. Sup. 6.51  $\pm$  2.64 0.00  $\pm$  0.00 23.72  $\pm$  8.75 30.89  $\pm$  7.21 13.31  $\pm$  4.42 26.45  $\pm$  3.74 6.57  $\pm$  3.01 14.38  $\pm$  6.09
REPLM (random all) Train 5.45  $\pm$  1.48 11.77  $\pm$  3.67 25.24  $\pm$  2.93 28.60  $\pm$  4.73 20.51  $\pm$  3.65 28.54  $\pm$  2.88 0.00  $\pm$  0.00 13.93  $\pm$  4.04
REPLM (random all) Dist. Sup. 5.64  $\pm$  1.87 11.47  $\pm$  3.78 23.76  $\pm$  3.43 24.73  $\pm$  4.62 14.47  $\pm$  3.10 24.09  $\pm$  3.17 5.48  $\pm$  2.89 13.65  $\pm$  3.79
REPLM (best $\ominus$) Train 0.00 25.81 39.42 36.70 17.20 32.43 13.33 8.00
REPLM (best $\ominus$) Dist. Sup. 23.33 21.51 40.85 33.85 14.12 26.92 10.00 10.53
REPLM (best $\oplus$) Train 0.00 24.18 39.42 39.63 17.82 38.46 10.67 6.67
REPLM (best $\oplus$) Dist. Sup. 22.22 29.21 37.24 30.81 14.81 31.58 11.43 10.53
Table 15: Ablation study. Comparing the performance across the in-context examples from distant supervision vs. human-annotated train set. Shown are F1 scores on each relation. (Part 6 / 12)
Method Context Source P172 P175 P176 P178 P179 P190 P194 P205
REPLM (random fixed) Train 13.59  $\pm$  5.05 36.38  $\pm$  8.61 11.89  $\pm$  5.31 22.27  $\pm$  3.21 14.80  $\pm$  3.95 0.00  $\pm$  0.00 12.67  $\pm$  3.37 20.50  $\pm$  7.77
REPLM (random fixed) Dist. Sup. 19.33  $\pm$  7.93 34.03  $\pm$  4.85 13.62  $\pm$  4.27 17.96  $\pm$  4.72 10.19  $\pm$  3.41 0.00  $\pm$  0.00 11.62  $\pm$  3.92 16.28  $\pm$  7.56
REPLM (random all) Train 11.62  $\pm$  3.91 38.49  $\pm$  2.49 12.14  $\pm$  2.80 22.09  $\pm$  2.07 12.95  $\pm$  1.74 0.00  $\pm$  0.00 13.66  $\pm$  4.93 22.61  $\pm$  7.69
REPLM (random all) Dist. Sup. 16.76  $\pm$  6.25 34.49  $\pm$  2.53 13.54  $\pm$  5.30 21.67  $\pm$  4.32 8.56  $\pm$  2.64 0.00  $\pm$  0.00 13.79  $\pm$  4.14 12.68  $\pm$  4.19
REPLM (best $\ominus$) Train 10.91 41.92 22.64 21.62 15.38 66.67 21.51 39.34
REPLM (best $\ominus$) Dist. Sup. 23.73 40.71 24.14 28.57 22.45 100.00 19.57 13.56
REPLM (best $\oplus$) Train 14.29 43.43 29.09 22.81 17.31 66.67 21.28 22.64
REPLM (best $\oplus$) Dist. Sup. 34.48 46.66 23.08 28.99 23.91 66.67 25.81 14.04
Table 16: Ablation study. Comparing the performance across the in-context examples from distant supervision vs. human-annotated train set. Shown are F1 scores on each relation. (Part 7 / 12)
Method Context Source P206 P241 P264 P272 P276 P279 P355 P361
REPLM (random fixed) Train 7.59  $\pm$  3.03 40.15  $\pm$  9.19 35.63  $\pm$  6.93 27.86  $\pm$  5.07 8.40  $\pm$  3.27 0.00  $\pm$  0.00 0.00  $\pm$  0.00 9.31  $\pm$  4.73
REPLM (random fixed) Dist. Sup. 6.68  $\pm$  3.53 36.97  $\pm$  9.98 28.62  $\pm$  9.75 28.67  $\pm$  3.68 8.32  $\pm$  3.07 0.00  $\pm$  0.00 14.43  $\pm$  6.32 7.36  $\pm$  4.70
REPLM (random all) Train 5.84  $\pm$  2.03 39.47  $\pm$  3.73 32.63  $\pm$  1.86 26.46  $\pm$  3.73 10.13  $\pm$  2.96 0.00  $\pm$  0.00 15.48  $\pm$  1.86 9.84  $\pm$  1.61
REPLM (random all) Dist. Sup. 0.00  $\pm$  0.00 33.24  $\pm$  3.41 29.66  $\pm$  2.15 28.10  $\pm$  4.05 9.88  $\pm$  2.77 0.00  $\pm$  0.00 11.26  $\pm$  5.08 7.66  $\pm$  1.82
REPLM (best $\ominus$) Train 11.57 46.75 32.29 24.56 17.27 6.56 21.74 18.93
REPLM (best $\ominus$) Dist. Sup. 14.63 43.24 30.77 35.09 20.00 11.11 30.77 27.12
REPLM (best $\oplus$) Train 8.20 46.58 29.91 21.43 20.44 2.94 17.78 23.10
REPLM (best $\oplus$) Dist. Sup. 18.18 45.07 28.35 37.29 22.76 10.91 33.33 25.66
Table 17: Ablation study. Comparing the performance across the in-context examples from distant supervision vs. human-annotated train set. Shown are F1 scores on each relation. (Part 8 / 12)
Method Context Source P364 P400 P403 P449 P463 P488 P495 P527
REPLM (random fixed) Train 24.96  $\pm$  6.35 36.17  $\pm$  10.65 22.47  $\pm$  3.76 29.08  $\pm$  6.31 19.08  $\pm$  2.69 9.85  $\pm$  2.21 15.27  $\pm$  2.74 11.89  $\pm$  2.03
REPLM (random fixed) Dist. Sup. 22.50  $\pm$  3.05 31.76  $\pm$  10.44 27.58  $\pm$  6.54 20.31  $\pm$  4.25 16.47  $\pm$  5.72 11.17  $\pm$  1.87 12.08  $\pm$  1.52 10.20  $\pm$  2.71
REPLM (random all) Train 21.75  $\pm$  5.54 41.02  $\pm$  4.49 21.09  $\pm$  3.61 28.86  $\pm$  3.35 16.58  $\pm$  4.92 12.66  $\pm$  4.46 12.95  $\pm$  1.50 10.42  $\pm$  2.31
REPLM (random all) Dist. Sup. 23.18  $\pm$  5.15 30.09  $\pm$  7.83 27.12  $\pm$  4.96 24.28  $\pm$  5.57 15.54  $\pm$  3.49 10.01  $\pm$  2.80 11.91  $\pm$  2.21 10.55  $\pm$  1.76
REPLM (best $\ominus$) Train 35.71 31.37 17.54 30.51 32.09 11.76 18.60 15.95
REPLM (best $\ominus$) Dist. Sup. 52.00 40.00 18.52 33.33 38.61 18.18 22.71 23.31
REPLM (best $\oplus$) Train 23.26 43.40 14.04 30.51 28.11 11.11 21.05 20.45
REPLM (best $\oplus$) Dist. Sup. 43.14 29.36 25.00 27.69 33.80 17.14 21.55 27.63
Table 18: Ablation study. Comparing the performance across the in-context examples from distant supervision vs. human-annotated train set. Shown are F1 scores on each relation. (Part 9 / 12)
Method Context Source P551 P569 P570 P571 P576 P577 P580 P582
REPLM (random fixed) Train 0.00  $\pm$  0.00 54.19  $\pm$  5.75 41.57  $\pm$  3.37 30.85  $\pm$  8.75 7.82  $\pm$  3.26 39.58  $\pm$  4.69 19.20  $\pm$  3.61 24.98  $\pm$  5.96
REPLM (random fixed) Dist. Sup. 29.09  $\pm$  8.91 51.90  $\pm$  8.21 44.07  $\pm$  3.16 24.60  $\pm$  11.18 8.85  $\pm$  2.46 32.29  $\pm$  5.03 18.35  $\pm$  5.51 0.00  $\pm$  0.00
REPLM (random all) Train 0.00  $\pm$  0.00 55.28  $\pm$  1.46 42.43  $\pm$  1.70 37.54  $\pm$  2.45 9.69  $\pm$  2.11 36.66  $\pm$  1.80 17.65  $\pm$  4.91 24.41  $\pm$  5.92
REPLM (random all) Dist. Sup. 0.00  $\pm$  0.00 55.43  $\pm$  1.43 39.35  $\pm$  1.82 23.17  $\pm$  2.44 5.74  $\pm$  2.03 36.61  $\pm$  2.18 19.23  $\pm$  3.65 11.71  $\pm$  4.98
REPLM (best $\ominus$) Train 18.18 57.10 44.35 39.73 9.23 43.08 25.81 35.90
REPLM (best $\ominus$) Dist. Sup. 36.36 60.95 46.92 38.96 14.49 46.78 27.40 31.11
REPLM (best $\oplus$) Train 18.18 56.56 45.55 36.49 5.88 42.28 22.95 34.15
REPLM (best $\oplus$) Dist. Sup. 36.36 62.25 46.96 40.27 20.00 45.35 18.46 28.57
Table 19: Ablation study. Comparing the performance across the in-context examples from distant supervision vs. human-annotated train set. Shown are F1 scores on each relation. (Part 10 / 12)
Method Context Source P585 P607 P674 P676 P706 P710 P737 P740
REPLM (random fixed) Train 16.10  $\pm$  4.22 16.02  $\pm$  3.99 25.00  $\pm$  5.89 54.95  $\pm$  6.90 0.00  $\pm$  0.00 8.42  $\pm$  6.01 0.00  $\pm$  0.00 23.41  $\pm$  4.25
REPLM (random fixed) Dist. Sup. 15.58  $\pm$  4.51 16.23  $\pm$  3.52 23.46  $\pm$  7.00 0.00  $\pm$  0.00 0.00  $\pm$  0.00 6.70  $\pm$  6.66 0.00  $\pm$  0.00 19.36  $\pm$  7.43
REPLM (random all) Train 18.76  $\pm$  4.51 16.61  $\pm$  1.99 22.46  $\pm$  2.71 58.90  $\pm$  2.15 11.93  $\pm$  2.47 9.23  $\pm$  5.93 0.00  $\pm$  0.00 22.03  $\pm$  8.04
REPLM (random all) Dist. Sup. 17.37  $\pm$  5.44 15.68  $\pm$  1.45 24.97  $\pm$  4.74 47.97  $\pm$  7.83 6.26  $\pm$  2.60 11.34  $\pm$  2.80 0.00  $\pm$  0.00 20.21  $\pm$  8.17
REPLM (best $\ominus$) Train 31.88 17.62 21.43 46.15 13.64 24.49 0.00 23.08
REPLM (best $\ominus$) Dist. Sup. 25.64 26.09 26.42 61.54 16.33 27.96 0.00 38.46
REPLM (best $\oplus$) Train 35.82 19.79 26.17 61.54 15.05 17.20 0.00 37.04
REPLM (best $\oplus$) Dist. Sup. 24.66 26.37 34.29 61.54 12.24 23.40 0.00 37.04
Table 20: Ablation study. Comparing the performance across the in-context examples from distant supervision vs. human-annotated train set. Shown are F1 scores on each relation. (Part 11 / 12)
Method Context Source P749 P800 P807 P840 P937 P1001 P1056 P1198
REPLM (random fixed) Train 11.47  $\pm$  2.41 20.02  $\pm$  4.19 0.00  $\pm$  0.00 35.43  $\pm$  6.25 23.90  $\pm$  4.76 13.51  $\pm$  5.41 0.00  $\pm$  0.00 0.00  $\pm$  0.00
REPLM (random fixed) Dist. Sup. 10.39  $\pm$  4.08 21.38  $\pm$  2.96 0.00  $\pm$  0.00 26.35  $\pm$  5.02 23.37  $\pm$  6.50 11.84  $\pm$  5.84 0.00  $\pm$  0.00 0.00  $\pm$  0.00
REPLM (random all) Train 11.32  $\pm$  2.73 20.02  $\pm$  2.55 0.00  $\pm$  0.00 35.41  $\pm$  8.47 22.45  $\pm$  5.03 12.60  $\pm$  1.68 0.00  $\pm$  0.00 0.00  $\pm$  0.00
REPLM (random all) Dist. Sup. 9.11  $\pm$  4.02 22.51  $\pm$  2.95 0.00  $\pm$  0.00 19.91  $\pm$  3.92 24.87  $\pm$  4.55 11.30  $\pm$  2.91 0.00  $\pm$  0.00 0.00  $\pm$  0.00
REPLM (best $\ominus$) Train 22.86 25.64 0.00 30.77 20.51 17.65 14.29 0.00
REPLM (best $\ominus$) Dist. Sup. 25.35 31.46 66.67 46.15 21.28 22.06 0.00 100.00
REPLM (best $\oplus$) Train 18.46 20.93 0.00 23.08 15.00 21.58 0.00 0.00
REPLM (best $\oplus$) Dist. Sup. 25.35 29.27 66.67 46.15 22.22 29.20 0.00 100.00
Table 21: Ablation study. Comparing the performance across the in-context examples from distant supervision vs. human-annotated train set. Shown are F1 scores on each relation. (Part 12 / 12)
Method Context Source P1336 P1344 P1365 P1366 P1376 P1412 P1441 P3373
REPLM (random fixed) Train 0.00  $\pm$  0.00 19.81  $\pm$  7.60 0.00  $\pm$  0.00 0.00  $\pm$  0.00 30.09  $\pm$  7.72 20.64  $\pm$  3.08 16.56  $\pm$  4.57 22.31  $\pm$  5.88
REPLM (random fixed) Dist. Sup. 0.00  $\pm$  0.00 20.41  $\pm$  6.54 0.00  $\pm$  0.00 0.00  $\pm$  0.00 34.86  $\pm$  10.12 24.65  $\pm$  4.02 15.52  $\pm$  5.30 22.74  $\pm$  5.03
REPLM (random all) Train 0.00  $\pm$  0.00 18.93  $\pm$  5.08 0.00  $\pm$  0.00 0.00  $\pm$  0.00 32.08  $\pm$  7.43 19.19  $\pm$  2.52 15.38  $\pm$  2.60 21.81  $\pm$  3.52
REPLM (random all) Dist. Sup. 0.00  $\pm$  0.00 16.99  $\pm$  3.11 0.00  $\pm$  0.00 0.00  $\pm$  0.00 33.58  $\pm$  3.41 25.84  $\pm$  5.77 14.65  $\pm$  4.25 25.44  $\pm$  4.98
REPLM (best $\ominus$) Train 0.00 33.33 0.00 0.00 29.27 17.39 16.97 24.04
REPLM (best $\ominus$) Dist. Sup. 31.58 45.54 0.00 0.00 52.38 42.35 20.00 20.11
REPLM (best $\oplus$) Train 0.00 33.01 0.00 0.00 34.15 33.33 18.60 30.37
REPLM (best $\oplus$) Dist. Sup. 31.58 43.40 0.00 0.00 68.29 34.48 29.71 16.48
Appendix E Implementation Details
We provide the details of our REPLM implementation in this section. We use GPT-JT [^13] https://huggingface.co/togethercomputer/GPT-JT-6B-v1 ($\sim$6B parameters) as our pre-trained LM for in-context few-shot learning. As the number of relations to be extracted is unknown in advance, we generate 200 tokens (for comparison, each extracted triplet consumes roughly 10-15 tokens) to ensure that our pre-trained LM can generate all relations it identifies.
As we use a fixed prefix for each extracted knowledge triplet at each line (e. g., “Relation:”), we easily identify if there is no further triplets extracted, simply from the absence of the prefix. We use a special separator token to easily parse the extracted subjects and objects. This separator is “<==>” in our experiments (which cannot be found in the original dataset and therefore cannot be confused with a natural text). We additionally inform our pre-trained LM about the task via starting our prompt with the instruction of the task. Here, we note that we have not done any prompt-tuning, since it is not the focus of this paper. For the output generation, we finally note that we use a greedy-decoding, e. g., not any sampling approach applied, which results in deterministic outputs given the input text. Example inputs and outputs can be found in Sec. G.
In our REPLM framework, we retrieve the semantically most-relevant in-context examples for each dev document. For this, we encode the documents via SBERT (Reimers and Gurevych, 2019) to calculate the embeddings and retrieve the most-relevant documents based on the cosine-similarity of the embeddings. We use the following fixed parameters in our framework (if not specified otherwise): $N=20$, $K=5$ $L=5$, $\tau=0.1$, and $\theta=0.2$. As the sentences in sentence-level relation extraction datasets (CONLL04, NYT, and ADE) are shorter than the documents in DocRED, we used more in-context examples for these datasets, which is $K=11$. The other parameters are the same as before.
We run all our experiments on NVIDIA Tesla V100-SXM2 32GB with a batch size 4. For DocRED, on average, each batch is processed in $\sim$17.80 seconds. As a result, the dev set (998 documents) is processed in $\sim$74.17 minutes for each relation type.
Appendix F Empirical Analysis of the Documents that Lack the Annotation
We conjecture that REPLM generates more comprehensive output than REBEL, not because REPLM the probability threshold $\theta$ is too low but because REPLM identifies annotations that are missing in the dataset. To validate this empirically, we thus adopt a simple yet effective strategy to predict if a given document contains information about the given relation. Fig. 5 plots the average cosine similarity between (i)  documents in the dev split and (ii) their top-$N$ nearest neighbors in distantly-supervised documents. The histogram shows two modes in distribution, where one corresponds to “known” relation and one to “missing” relations. In the overlap of two distributions, there are semantically similar documents that potentially include the relation but lack the annotation.
(a) P19
(b) P27
(c) P131
(d) P175
(e) P570
(f) P577
Figure 5: Histogram of average cosine similarity between documents and their top-$N$ neighbors for two example relation types.
Appendix G Example Prompts and Outputs
In the following, we provide examples for different prompts as input (in red) and the corresponding output (in blue).
G.1 P17 (country)
Input Prompt:
Your task is to identify all the unique knowledge triplets of ’country’ for a given context. Knowledge triplet will be ordered as relation, subject, and object, which are separated by <==>. If there are multiple triplets, list each of them in a new line. Follow the example context-relation pairs for the formatting of your output.
Context: IBM Laboratory Vienna was an IBM research laboratory based in Vienna , Austria . The laboratory started with a group led by Heinz Zemanek that moved from the Technische Hochschule ( now the Technical University of Vienna ) . Initially , the group worked on computer hardware projects . Later a compiler for the ALGOL 60 programming language was produced . The group built on ideas of Calvin C. Elgot , Peter Landin , and John McCarthy , to create an operational semantics that could define the whole of IBM ’s PL / I programming language . The meta - language used for this was dubbed by people outside the laboratory as the Vienna Definition Language ( VDL ) . These descriptions were used for compiler design research into compiler design during 1968 - 70 . The formal method VDM ( Vienna Development Method ) was a result of research at the laboratory by Dines Bjørner , Cliff Jones , Peter Lucas , and others .
Relation: (country <==> Vienna <==> Austria)
Relation: (country <==> Technical University of Vienna <==> Austria)
Context: The School of Engineering of Juiz de Fora ( ) was an engineering college in the city of Juiz de Fora , Brazil . It is now the engineering faculty of the Federal University of Juiz de Fora ( UFJF ) . The former president of Brazil Itamar Franco was an alumnus . It was set up in 1914 in the city of Juiz de Fora , Minas Gerais state , Brazil , and taught a five - year course of Civil and Eletrotechnic Engineering . In 1960 , the school joined the Medicine , Pharmacy and Law schools of that city to found the UFJF . Nowadays , the Faculty of Engineering provides courses in civil , production , electrical ( divided into telecommunication , energy , power , electronic , robotic and automation systems ) , mechanical , computer , sanitary and environmental engineering , and architecture . Relation: (country <==> School of Engineering of Juiz de Fora <==> Brazil)
Relation: (country <==> Juiz de Fora <==> Brazil)
Relation: (country <==> UFJF <==> Brazil)
Relation: (country <==> Minas Gerais <==> Brazil)
Context: Bizrate Insights Inc. , doing business as Bizrate Insights , is a market research company , providing consumer ratings information to over 6,000 retailers and publishers across the United States , United Kingdom , France , Germany , and Canada . Bizrate Insights is a Meredith Corporation company based in Los Angeles , CA . Bizrate Insights provides services to both businesses and consumers in two different ways : consumers have access to ratings and reviews from verified buyers that help to inform their purchase decisions . This feedback can be found on the Bizrate website and is syndicated across the web to major search engines such as Google and Bing . Bizrate Insights provides businesses with customer satisfaction insights about consumers , advanced analytics , and competitive benchmarks across all types of online retail industries . Bizrate Insights also provides industry research to analysts at Forrester Research and Internet Retailer for publication and studies . Relation: (country <==> Meredith Corporation <==> the United States)
Relation: (country <==> Los Angeles <==> the United States)
Relation: (country <==> Google <==> the United States)
Relation: (country <==> Bing <==> the United States)
Context: The Universidade Positivo ( abbreviated UP ) is a private universities of the State of Paraná , Brazil . Universidade Positivo ’s campus is in the Campo Comprido district of Curitiba and occupies an area of 420,004 m2 . Its 114,000 m2 of installations . Universidade Positivo currently offers 27 undergraduate programs , a doctoral program , three master ’s degree programs , certificate of graduate study programs and extension courses . The 6,300 m2 library can hold 864 users at a time . Its collection consists of nearly 115,000 volumes . In a special room the library houses the personal collection of the late cabinet minister , diplomat and professor , Roberto de Oliveira Campos . It includes diplomas , medals and awards , in addition to its 8,426 volumes .
Relation: (country <==> Universidade Positivo <==> Brazil)
Relation: (country <==> State of Paraná <==> Brazil)
Relation: (country <==> Curitiba <==> Brazil)
Context: Softtek is a Mexican - based information technology company , operating in North America , Latin America , Europe and Asia . As of 2010 it was the largest private IT vendor in Latin America . Headquartered in Monterrey , Mexico , the company has 12,000 associates in Mexico and abroad . The company offers application software development , testing , security and support ; business process outsourcing ( BPO ) ; and IT infrastructure management , security and support to more than 400 corporations in more than 20 countries . It also acts as a value added reseller ( VAR ) for SAP SE , Informatica , Cognos , Business Objects and other software products . The company has trademarked the term " nearshoring " to describe the provision of outsourced services to customers in other countries that are in proximity .
Relation: (country <==> Softtek <==> Mexico)
Relation: (country <==> Monterrey <==> Mexico)
Context: IBM Research – Brazil is one of twelve research laboratories comprising IBM Research , its first in South America . It was established in June 2010 , with locations in São Paulo and Rio de Janeiro . Research focuses on Industrial Technology and Science , Systems of Engagement and Insight , Social Data Analytics and Natural Resources Solutions . The new lab , IBM ’s ninth at the time of opening and first in 12 years , underscores the growing importance of emerging markets and the globalization of innovation . In collaboration with Brazil ’s government , it will help IBM to develop technology systems around natural resource development and large - scale events such as the 2016 Summer Olympics . Engineer and associate lab director Ulisses Mello explains that IBM has four priority areas in Brazil : " The main area is related to natural resources management , involving oil and gas , mining and agricultural sectors . The second is the social data analytics segment that comprises the analysis of data generated from social networking sites [ such as Twitter or Facebook ] , which can be applied , for example , to financial analysis . The third strategic area is nanotechnology applied to the development of the smarter devices for the intermittent production industry . This technology can be applied to , for example , blood testing or recovering oil from existing fields . And the last one is smarter cities .
Output:
Relation: (country <==> IBM Research – Brazil <==> Brazil)
Relation: (country <==> São Paulo <==> Brazil)
Relation: (country <==> Rio de Janeiro <==> Brazil)
Relation: (country <==> 2016 Summer Olympics <==> Brazil)
Relation: (country <==> Industrial Technology and Science <==> Brazil)
Correctness:
For relation P17 (country), the given dev document is annotated with two knowledge triplets (country, São Paulo, Brazil) and (country, Rio de Janeiro, Brazil). For this document, we find that the annotators did not consider the country information of the event “2016 Summer Olympics” and the organization “IBM Research – Brazil” for their manual annotation. Validating from an external knowledge base, we observe that our REPLM framework identifies two more knowledge triplets correctly, (country, 2016 Summer Olympics, Brazil) and (country, Industrial Technology and Science, Brazil).
G.2 P118 (league)
Input Prompt:
Your task is to identify all the unique knowledge triplets of ’league’ for a given context. Knowledge triplet will be ordered as relation, subject, and object, which are separated by <==>. If there are multiple triplets, list each of them in a new line. Follow the example context-relation pairs for the formatting of your output.
Context: Richard Ian Wright ( born 5 November 1977 ) is an English retired professional footballer who played as a goalkeeper . He currently works as a coach for Manchester City . Born in Ipswich , he joined Ipswich Town as a trainee , going on to play for the club 298 times between 1995 and 2001 . He then moved to Premier League club Arsenal , before being signed by Everton in 2002 , where he spent five years . A brief spell on loan from West Ham United with Southampton was followed by a transfer back to Ipswich Town . After a short spell at Sheffield United , a third stint at Ipswich and a brief time at Preston North End , he joined Premier League champions Manchester City on a free transfer in 2012 . After four years at City , during which he did not play at all , he announced his retirement in May 2016 . He remained with City as a coach under new manager Pep Guardiola . He was a member of the England squad , earning two caps , and was included in their squad for UEFA Euro 2000 .
Relation: (league <==> Manchester City <==> Premier League)
Relation: (league <==> Arsenal <==> Premier League)
Relation: (league <==> Everton <==> Premier League)
Relation: (league <==> West Ham United <==> Premier League)
Relation: (league <==> Southampton <==> Premier League)
Context: Ashley Renaldo Chambers ( born 1 March 1990 ) is an English professional footballer who plays as a winger or a striker for club Kidderminster Harriers . Chambers started his career with Leicester City , making his first - team debut in 2005 at the age of 15 in a League Cup match against Blackpool , which made him the youngest player in the club ’s history . In 2009 , he joined League One club Wycombe Wanderers on loan . This was followed by a loan period with League Two club Grimsby Town . He signed for Conference Premier club York City on loan in November 2010 before signing permanently . He won in the 2012 FA Trophy Final and 2012 Conference Premier play - off Final with York at Wembley Stadium , the latter seeing the club promoted into League Two . Chambers joined Conference Premier club Cambridge United in 2014 .
Relation: (league <==> Kidderminster Harriers <==> Conference Premier)
Relation: (league <==> Leicester City <==> League One)
Context: Roy Eric Carroll ( born 30 September 1977 ) is a Northern Irish professional footballer who plays as a goalkeeper for NIFL Premiership side Linfield . He is best known for his spells at Wigan Athletic , Manchester United ( where he won a Premier League winners medal and the 2004 FA Cup ) and Olympiacos ( where he won the Greek Superleague three times and the Greek Cup twice ) . He has also represented Northern Ireland 45 times at full international level , gaining his first cap in 1997 , aged 19 . Carroll has also had a one - game managerial career , leading Barnet to a 2 - 1 victory in the 2011 Herts Senior Cup final against Stevenage . Therefore , Carroll holds the unusual honour of having won a trophy in his only game as a manager .
Relation: (league <==> Linfield <==> NIFL Premiership)
Relation: (league <==> Manchester United <==> Premier League)
Relation: (league <==> Olympiacos <==> Greek Superleague)
Context: Brian Christopher Deane ( born 7 February 1968 ) is an English football coach and former player whose most recent position was as the manager of the Norwegian side Sarpsborg 08 . During his playing career , he played as forward from 1985 until 2006 . He was the scorer of the first ever goal in the FA Premier League in 1992 , when he was a Sheffield United player . Deane also played in the Premier League for Leeds United and Middlesbrough as well as playing top - flight football in Portugal and Australia for Benfica and Perth Glory respectively . He also played in The Football League for Doncaster Rovers , Leicester City , West Ham United and Sunderland before finishing his playing career in 2006 with a brief spell back at Sheffield United . Deane was capped three times by England .
Relation: (league <==> Middlesbrough <==> FA Premier League)
Relation: (league <==> Leicester City <==> FA Premier League)
Relation: (league <==> West Ham United <==> FA Premier League)
Context: John Stones ( born 28 May 1994 ) is an English professional footballer who plays for club Manchester City and the English national team . Mainly a centre back , he can also play as a right back . Stones began his career with Barnsley , making his first - team debut in the Championship in March 2012 as a 17-year - old . He joined Premier League club Everton for around £ 3 million in January 2013 and amassed 95 appearances over four seasons . In August 2016 , he signed for Manchester City for an initial £ 47.5 million with add - ons . He won the Premier League and EFL Cup in 2018 . Stones made his senior debut for England in May 2014 after previously being capped by England youth teams at under-19 , under-20 and under-21 levels . He was chosen in England ’s squads for UEFA Euro 2016 and the 2018 FIFA World Cup .
Relation: (league <==> Manchester City <==> Premier League)
Relation: (league <==> Manchester City <==> EFL Cup)
Relation: (league <==> Barnsley <==> Championship)
Relation: (league <==> Everton <==> Premier League)
Context: Andrew Alexander Cole ( born 15 October 1971 ) is an English former professional footballer . Playing as a striker , his career lasted from 1988 to 2008 . He is most notably remembered for his time in the Premier League , with Manchester United , where he spent six years of his career , winning numerous trophies in the process . He also played in the top division of English football for Arsenal , Newcastle United , Blackburn Rovers , Fulham , Manchester City , Portsmouth and Sunderland , as well as in the Football League for Bristol City , Birmingham City , Burnley and Nottingham Forest . He is the third - highest goalscorer in Premier League history with 187 goals . Cole has the distinction of being one of the few players in England to have swept all possible honours in the English game , including the PFA Young Player of the Year award , as well as the coveted UEFA Champions League title . Cole was also capped 15 times for the England national team between 1995 and 2001 , scoring once against Albania in a 2002 FIFA World Cup qualifier .
Output:
Relation: (league <==> Manchester United <==> Premier League)
Relation: (league <==> Newcastle United <==> Premier League)
Relation: (league <==> Blackburn Rovers <==> Premier League)
Relation: (league <==> Fulham <==> Premier League)
Relation: (league <==> Manchester City <==> Premier League)
Relation: (league <==> Portsmouth <==> Premier League)
Relation: (league <==> Sunderland <==> Premier League)
Correctness:
For relation P118 (league), the given dev document is annotated with only one knowledge triplet (league, Manchester United, Premier League). In this example, our pre-trained LM identifies 6 more knowledge triplets via in-context few-shot learning. At the time of writing, the following triplets are also correct (league, Newcastle United, Premier League), (league, Fulham, Premier League), (league, Manchester City, Premier League). We note that although such information is outdated, Blackburn Rovers, Portsmouth, and Sunderland can be inferred as in Premier League from the same passage. This also highlights the need that the relation extraction methods should additionally infer (if possible) for which time period the extracted information is valid.
G.3 P272 (production company)
Input Prompt:
Your task is to identify all the unique knowledge triplets of ’production company’ for a given context. Knowledge triplet will be ordered as relation, subject, and object, which are separated by <==>. If there are multiple triplets, list each of them in a new line. Follow the example context-relation pairs for the formatting of your output.
Context: Ron Moody ( born Ronald Moodnick , 8 January 1924 - 11 June 2015 ) was an English actor , singer , composer and writer best known for his portrayal of Fagin in Oliver ! ( 1968 ) and its 1983 Broadway revival . Moody earned a Golden Globe Award and an Academy Award nomination for the film , as well as a Tony Award nomination for the stage production . Other notable projects include The Mouse on the Moon ( 1963 ) , Mel Brooks ’ The Twelve Chairs ( 1970 ) and Flight of the Doves ( 1971 ) , in which Moody shared the screen with Oliver ! co - star Jack Wild . Moody holds the peculiar distinction of having portrayed the wizard Merlin in two Disney films , Unidentified Flying Oddball ( 1979 ) and A Kid in King Arthur ’s Court ( 1995 ) .
Relation: (production company <==> Unidentified Flying Oddball <==> Disney)
Context: The Beastmaster is a 1982 sword and sorcery film directed by Don Coscarelli and starring Marc Singer , Tanya Roberts , John Amos and Rip Torn loosely based on the novel The Beast Master by Andre Norton . The film is about a child who is stolen from his mother ’s womb by a witch . The child grows into Dar , who has the ability to communicate telepathically with animals . Dar grows up in a village where he learns to do battle . But the village is destroyed by a race of beast - like warriors under the control of the sorcerer Maax . Dar vows revenge and travels with new friends to stop Maax from causing any more problems . Commercially The Beastmaster was not considered a box office success during its original cinematic run ; however later it received extensive television exposure and success on cable in the American market on channels TBS and HBO . The original film spawned two sequels as well as a syndicated television series that chronicled the further adventures of Dar .
Relation: (production company <==> Beastmaster <==> Don Coscarelli)
Context: Monkeybone is a 2001 American black comedy dark fantasy film directed by Henry Selick , written by Sam Hamm , and produced by Selick , Hamm , Mark Radcliffe , Michael Barnathan , and Chris Columbus . The film combines live - action with stop - motion animation . Based on Kaja Blackley ’s graphic novel Dark Town , the film stars an ensemble cast led by Brendan Fraser , Bridget Fonda , and Whoopi Goldberg with Rose McGowan , Dave Foley , Giancarlo Esposito , Megan Mullally , Lisa Zane , Chris Kattan , John Turturro , and an uncredited Thomas Haden Church . Theatrically released on February 23 , 2001 by 20th Century Fox , the film was a box office bomb and received generally negative critical reviews .
Relation: (production company <==> Monkeybone <==> 20th Century Fox)
Context: TaleSpin is an American animated television series based in the fictional city of Cape Suzette , that first aired in 1990 as a preview on The Disney Channel and later that year as part of The Disney Afternoon , with characters adapted from Disney ’s 1967 animated feature The Jungle Book , which was theatrically rereleased in the summer before this show premiered in the fall . The name of the show is a play on tailspin , the rapid descent of an aircraft in a steep spiral . The two words in the show ’s name , tale and spin , are a way to describe telling a story . The show is one of ten Disney Afternoon shows to use established Disney characters as the main characters , with the others being Darkwing Duck , DuckTales , Chip ’ n Dale Rescue Rangers , Goof Troop , Bonkers , Quack Pack , Aladdin , Timon & Pumbaa and Jungle Cubs . It is also one of the two animated television series based on The Jungle Book along with Jungle Cubs .
Relation: (production company <==> Aladdin <==> Disney)
Context: Tarzan ’s Hidden Jungle is a 1955 black - and - white film from RKO Pictures directed by Harold D. Schuster and starring Gordon Scott in his first film as Tarzan , taking over the role from Lex Barker , who had in turn followed Johnny Weissmuller in the series . The film about Edgar Rice Burroughs ’ ape - man also features Vera Miles and Jack Elam . It was the last of twelve Tarzan pictures released by RKO . Tarzan ’s mate , Jane , does not appear in the film . Tarzan at first seems to show more than casual interest in Miles ’ character , but ultimately there is no romance . In real life , Scott and Miles were married after the film was completed . Scott eventually played Tarzan in six movies over a five - year span .
Relation: (production company <==> Tarzan ’s Hidden Jungle <==> RKO Pictures)
Context: King Louie is a fictional character introduced in Walt Disney ’s 1967 animated musical film , The Jungle Book . Unlike the majority of the adapted characters in the film , Louie was not featured in Rudyard Kipling ’s original works . King Louie was portrayed as an orangutan who was the leader of the other jungle primates , and who attempted to gain knowledge of fire from Mowgli , in order to become more human . King Louie was voiced by Louis Prima in the original 1967 film . Initially , the filmmakers considered Louis Armstrong for the role , but to avoid the likely controversy that would result from casting a black person to voice an ape , they instead chose Prima , a white singer . Christopher Walken voiced King Louie in the 2016 live - action / CGI The Jungle Book film .
Output:
Relation: (production company <==> The Jungle Book <==> Walt Disney)
Correctness:
For relation P637 (production company), there was no annotation for the given dev document. However, our framework identifies the knowledge triplet (production company, The Jungle Book, Walt Disney). This can easily be inferred from the first sentence of the document.
Appendix H Full Results on Overall Performance
Tables 22 to 33 provide the detailed results of relation extraction of all methods on all relation types. For methods with random in-context examples, the performance may be subject to variability across which seed is picked (whereas the performance is deterministic for the other methods), and, hence, we report the standard deviation for this subset of the methods by averaging the performance across 10 runs. The evaluation is done based on the human annotations of the dev documents.
When we compare the different variants of our framework, the results confirm our choice in the complete REPLM framework. Specifically, we find that (1) retrieving the best in-context examples improves the performance compared to random examples (i. e., REPLM (best $\ominus$) and REPLM (best $\oplus$) vs. REPLM (random fixed) and REPLM (random all)) and (2) our complete REPLM framework brings a significant improvement over REPLM (best $\ominus$) and REPLM (best $\oplus$) by aggregating multiple sets of most relevant in-context examples, thus establishing the importance of using multiple sets. We further compare our framework against the state-of-the-art relation extraction method REBEL and find that, overall, our REPLM and REPLM (params adj) perform the best.
Table 22: Full results in overall performance (Section 6.1). Shown are F1 scores on each relation. (Part 1 / 12)
Method P6 P17 P19 P20 P22 P25 P26 P27
REBEL 0.00 20.66 50.89 29.41 40.86 38.10 41.32 5.69
REBEL-sent 4.08 25.73 45.95 29.85 33.71 30.00 41.84 11.37
REPLM (random fixed) 24.28  $\pm$  10.53 11.34  $\pm$  4.48 66.67  $\pm$  12.71 49.80  $\pm$  16.07 9.76  $\pm$  4.90 0.00  $\pm$  0.00 26.73  $\pm$  4.38 22.41  $\pm$  2.53
REPLM (random all) 25.90  $\pm$  2.44 11.05  $\pm$  0.46 68.49  $\pm$  3.09 48.35  $\pm$  4.82 9.51  $\pm$  2.86 0.00  $\pm$  0.00 26.74  $\pm$  3.83 23.26  $\pm$  1.39
REPLM (best $\ominus$) 35.96 24.60 71.38 50.39 18.18 15.38 33.33 29.14
REPLM (best $\oplus$) 35.96 24.02 68.63 62.50 17.70 7.41 22.97 28.33
REPLM 34.78 27.74 59.06 40.41 17.17 28.00 37.17 30.71
REPLM (params adj) 29.79 27.85 77.49 59.77 15.54 22.54 40.24 34.64
Table 23: Full results in overall performance (Section 6.1). Shown are F1 scores on each relation. (Part 2 / 12)
Method P30 P31 P35 P36 P37 P39 P40 P50
REBEL 6.40 9.84 0.00 57.89 0.00 46.15 39.02 42.76
REBEL-sent 10.29 16.16 3.70 55.81 0.00 33.33 35.00 51.81
REPLM (random fixed) 13.58  $\pm$  1.66 5.21  $\pm$  2.48 28.79  $\pm$  4.96 18.67  $\pm$  5.96 22.59  $\pm$  8.79 0.00  $\pm$  0.00 12.97  $\pm$  3.62 27.66  $\pm$  4.05
REPLM (random all) 14.75  $\pm$  2.18 6.33  $\pm$  2.15 26.96  $\pm$  3.17 14.85  $\pm$  5.65 24.13  $\pm$  5.48 0.00  $\pm$  0.00 11.63  $\pm$  3.12 26.74  $\pm$  3.45
REPLM (best $\ominus$) 31.25 6.59 31.46 47.06 31.17 25.00 20.38 38.60
REPLM (best $\oplus$) 30.85 6.82 37.21 45.28 22.78 25.00 26.42 37.66
REPLM 39.11 14.29 32.56 52.46 27.91 8.70 17.56 41.12
REPLM (params adj) 38.94 14.04 29.58 52.83 29.73 12.50 20.38 41.12
Table 24: Full results in overall performance (Section 6.1). Shown are F1 scores on each relation. (Part 3 / 12)
Method P54 P57 P58 P69 P86 P102 P108 P112
REBEL 38.06 48.53 19.05 41.38 26.67 49.32 26.87 27.78
REBEL-sent 25.22 40.60 5.26 5.31 16.90 60.61 26.09 20.51
REPLM (random fixed) 42.71  $\pm$  10.02 34.55  $\pm$  4.59 27.65  $\pm$  6.61 53.00  $\pm$  9.62 16.16  $\pm$  8.57 30.82  $\pm$  9.53 33.29  $\pm$  4.49 14.30  $\pm$  8.81
REPLM (random all) 43.80  $\pm$  3.48 32.63  $\pm$  3.59 27.32  $\pm$  3.10 57.18  $\pm$  3.74 13.23  $\pm$  4.60 32.85  $\pm$  3.32 32.67  $\pm$  5.06 16.06  $\pm$  6.68
REPLM (best $\ominus$) 48.30 47.62 34.38 57.47 23.26 44.44 30.19 32.65
REPLM (best $\oplus$) 40.93 43.43 39.34 58.29 40.45 40.48 34.29 34.78
REPLM 45.78 43.70 30.77 56.25 51.33 42.26 31.88 31.43
REPLM (params adj) 43.42 45.99 35.82 64.41 51.33 45.90 34.00 37.21
Table 25: Full results in overall performance (Section 6.1). Shown are F1 scores on each relation. (Part 4 / 12)
Method P118 P123 P127 P131 P136 P137 P140 P150
REBEL 34.21 27.59 13.48 31.01 19.05 16.67 4.65 33.49
REBEL-sent 40.74 30.11 18.69 28.61 16.67 8.16 14.81 39.53
REPLM (random fixed) 32.16  $\pm$  5.80 19.81  $\pm$  4.20 11.10  $\pm$  4.41 14.32  $\pm$  2.81 19.70  $\pm$  6.09 9.99  $\pm$  3.84 0.00  $\pm$  0.00 23.36  $\pm$  2.23
REPLM (random all) 33.51  $\pm$  5.47 19.87  $\pm$  3.15 10.50  $\pm$  3.94 15.54  $\pm$  0.65 21.93  $\pm$  6.65 9.83  $\pm$  2.77 0.00  $\pm$  0.00 21.91  $\pm$  1.52
REPLM (best $\ominus$) 44.04 30.19 18.49 25.50 20.00 12.90 14.17 31.28
REPLM (best $\oplus$) 37.84 19.82 18.03 26.59 26.09 9.84 13.53 30.25
REPLM 39.75 32.47 21.52 29.89 18.18 17.72 14.29 34.36
REPLM (params adj) 46.28 35.77 21.79 29.79 18.18 17.95 14.55 33.75
Table 26: Full results in overall performance (Section 6.1). Shown are F1 scores on each relation. (Part 5 / 12)
Method P155 P156 P159 P161 P162 P166 P170 P171
REBEL 2.70 7.02 22.02 37.96 10.91 28.92 28.07 16.00
REBEL-sent 16.33 24.39 35.90 17.74 7.27 38.64 31.58 13.79
REPLM (random fixed) 6.51  $\pm$  2.64 0.00  $\pm$  0.00 23.72  $\pm$  8.75 30.89  $\pm$  7.21 13.31  $\pm$  4.42 26.45  $\pm$  3.74 6.57  $\pm$  3.01 14.38  $\pm$  6.09
REPLM (random all) 5.64  $\pm$  1.87 11.47  $\pm$  3.78 23.76  $\pm$  3.43 24.73  $\pm$  4.62 14.47  $\pm$  3.10 24.09  $\pm$  3.17 5.48  $\pm$  2.89 13.65  $\pm$  3.79
REPLM (best $\ominus$) 23.33 21.51 40.85 33.85 14.12 26.92 10.00 10.53
REPLM (best $\oplus$) 22.22 29.21 37.24 30.81 14.81 31.58 11.43 10.53
REPLM 16.09 29.93 45.03 45.23 23.81 26.42 15.52 14.63
REPLM (params adj) 16.47 30.07 42.11 46.21 23.64 27.59 17.50 10.81
Table 27: Full results in overall performance (Section 6.1). Shown are F1 scores on each relation. (Part 6 / 12)
Method P172 P175 P176 P178 P179 P190 P194 P205
REBEL 0.00 47.90 13.04 28.32 0.00 0.00 10.00 11.76
REBEL-sent 0.00 45.88 32.79 28.04 0.00 0.00 17.39 16.22
REPLM (random fixed) 19.33  $\pm$  7.93 34.03  $\pm$  4.85 13.62  $\pm$  4.27 17.96  $\pm$  4.72 10.19  $\pm$  3.41 0.00  $\pm$  0.00 11.62  $\pm$  3.92 16.28  $\pm$  7.56
REPLM (random all) 16.76  $\pm$  6.25 34.49  $\pm$  2.53 13.54  $\pm$  5.30 21.67  $\pm$  4.32 8.56  $\pm$  2.64 0.00  $\pm$  0.00 13.79  $\pm$  4.14 12.68  $\pm$  4.19
REPLM (best $\ominus$) 23.73 40.71 24.14 28.57 22.45 100.00 19.57 13.56
REPLM (best $\oplus$) 34.48 46.66 23.08 28.99 23.91 66.67 25.81 14.04
REPLM 32.91 52.19 28.17 31.75 27.27 66.67 30.65 25.35
REPLM (params adj) 30.77 53.27 29.73 31.75 23.30 66.67 21.51 27.12
Table 28: Full results in overall performance (Section 6.1). Shown are F1 scores on each relation. (Part 7 / 12)
Method P206 P241 P264 P272 P276 P279 P355 P361
REBEL 34.11 26.42 42.61 44.90 19.05 4.55 6.06 37.45
REBEL-sent 31.25 12.24 27.39 48.00 26.42 15.15 23.53 26.09
REPLM (random fixed) 6.68  $\pm$  3.53 36.97  $\pm$  9.98 28.62  $\pm$  9.75 28.67  $\pm$  3.68 8.32  $\pm$  3.07 0.00  $\pm$  0.00 14.43  $\pm$  6.32 7.36  $\pm$  4.70
REPLM (random all) 0.00  $\pm$  0.00 33.24  $\pm$  3.41 29.66  $\pm$  2.15 28.10  $\pm$  4.05 9.88  $\pm$  2.77 0.00  $\pm$  0.00 11.26  $\pm$  5.08 7.66  $\pm$  1.82
REPLM (best $\ominus$) 14.63 43.24 30.77 35.09 20.00 11.11 30.77 27.12
REPLM (best $\oplus$) 18.18 45.07 28.35 37.29 22.76 10.91 33.33 25.66
REPLM 20.51 46.00 45.54 36.36 30.12 9.09 22.22 25.39
REPLM (params adj) 11.57 53.12 46.75 36.36 32.79 8.96 28.07 25.33
Table 29: Full results in overall performance (Section 6.1). Shown are F1 scores on each relation. (Part 8 / 12)
Method P364 P400 P403 P449 P463 P488 P495 P527
REBEL 0.00 36.54 38.46 0.00 43.90 17.39 3.54 37.04
REBEL-sent 0.00 36.17 44.12 0.00 40.23 8.33 5.69 34.45
REPLM (random fixed) 22.50  $\pm$  3.05 31.76  $\pm$  10.44 27.58  $\pm$  6.54 20.31  $\pm$  4.25 16.47  $\pm$  5.72 11.17  $\pm$  1.87 12.08  $\pm$  1.52 10.20  $\pm$  2.71
REPLM (random all) 23.18  $\pm$  5.15 30.09  $\pm$  7.83 27.12  $\pm$  4.96 24.28  $\pm$  5.57 15.54  $\pm$  3.49 10.01  $\pm$  2.80 11.91  $\pm$  2.21 10.55  $\pm$  1.76
REPLM (best $\ominus$) 52.00 40.00 18.52 33.33 38.61 18.18 22.71 23.31
REPLM (best $\oplus$) 43.14 29.36 25.00 27.69 33.80 17.14 21.55 27.63
REPLM 39.34 34.01 27.03 38.10 42.07 14.81 22.41 21.74
REPLM (params adj) 38.10 35.21 21.28 39.39 41.05 20.00 21.82 23.49
Table 30: Full results in overall performance (Section 6.1). Shown are F1 scores on each relation. (Part 9 / 12)
Method P551 P569 P570 P571 P576 P577 P580 P582
REBEL 0.00 50.51 39.19 50.93 0.00 35.32 21.05 0.00
REBEL-sent 25.00 51.32 44.44 42.59 0.00 42.79 23.81 0.00
REPLM (random fixed) 29.09  $\pm$  8.91 51.90  $\pm$  8.21 44.07  $\pm$  3.16 24.60  $\pm$  11.18 8.85  $\pm$  2.46 32.29  $\pm$  5.03 18.35  $\pm$  5.51 0.00  $\pm$  0.00
REPLM (random all) 0.00  $\pm$  0.00 55.43  $\pm$  1.43 39.35  $\pm$  1.82 23.17  $\pm$  2.44 5.74  $\pm$  2.03 36.61  $\pm$  2.18 19.23  $\pm$  3.65 11.71  $\pm$  4.98
REPLM (best $\ominus$) 36.36 60.95 46.92 38.96 14.49 46.78 27.40 31.11
REPLM (best $\oplus$) 36.36 62.25 46.96 40.27 20.00 45.35 18.46 28.57
REPLM 28.57 55.73 44.56 39.71 17.78 52.39 29.17 43.14
REPLM (params adj) 33.33 61.29 50.66 46.78 19.35 54.40 35.48 46.15
Table 31: Full results in overall performance (Section 6.1). Shown are F1 scores on each relation. (Part 10 / 12)
Method P585 P607 P674 P676 P706 P710 P737 P740
REBEL 11.32 42.17 28.26 0.00 5.88 25.32 0.00 12.50
REBEL-sent 28.57 38.64 21.15 0.00 2.94 20.22 0.00 10.53
REPLM (random fixed) 15.58  $\pm$  4.51 16.23  $\pm$  3.52 23.46  $\pm$  7.00 0.00  $\pm$  0.00 0.00  $\pm$  0.00 6.70  $\pm$  6.66 0.00  $\pm$  0.00 19.36  $\pm$  7.43
REPLM (random all) 17.37  $\pm$  5.44 15.68  $\pm$  1.45 24.97  $\pm$  4.74 47.97  $\pm$  7.83 6.26  $\pm$  2.60 11.34  $\pm$  2.80 0.00  $\pm$  0.00 20.21  $\pm$  8.17
REPLM (best $\ominus$) 25.64 26.09 26.42 61.54 16.33 27.96 0.00 38.46
REPLM (best $\oplus$) 24.66 26.37 34.29 61.54 12.24 23.40 0.00 37.04
REPLM 30.11 33.20 31.25 47.06 23.02 30.19 0.00 31.58
REPLM (params adj) 28.26 27.69 23.30 46.15 17.39 23.66 0.00 31.25
Table 32: Full results in overall performance (Section 6.1). Shown are F1 scores on each relation. (Part 11 / 12)
Method P749 P800 P807 P840 P937 P1001 P1056 P1198
REBEL 4.26 39.47 0.00 0.00 0.00 15.05 0.00 0.00
REBEL-sent 32.35 43.90 0.00 0.00 0.00 1.94 0.00 0.00
REPLM (random fixed) 10.39  $\pm$  4.08 21.38  $\pm$  2.96 0.00  $\pm$  0.00 26.35  $\pm$  5.02 23.37  $\pm$  6.50 11.84  $\pm$  5.84 0.00  $\pm$  0.00 0.00  $\pm$  0.00
REPLM (random all) 9.11  $\pm$  4.02 22.51  $\pm$  2.95 0.00  $\pm$  0.00 19.91  $\pm$  3.92 24.87  $\pm$  4.55 11.30  $\pm$  2.91 0.00  $\pm$  0.00 0.00  $\pm$  0.00
REPLM (best $\ominus$) 25.35 31.46 66.67 46.15 21.28 22.06 0.00 100.00
REPLM (best $\oplus$) 25.35 29.27 66.67 46.15 22.22 29.20 0.00 100.00
REPLM 24.49 29.27 50.00 34.29 28.57 28.57 0.00 100.00
REPLM (params adj) 21.05 32.56 50.00 30.77 21.74 28.57 0.00 66.67
Table 33: Full results in overall performance (Section 6.1). Shown are F1 scores on each relation. (Part 12 / 12)
Method P1336 P1344 P1365 P1366 P1376 P1412 P1441 P3373
REBEL 0.00 0.00 18.18 18.18 25.00 0.00 27.78 41.81
REBEL-sent 0.00 0.00 33.33 23.53 24.00 0.00 18.54 46.35
REPLM (random fixed) 0.00  $\pm$  0.00 20.41  $\pm$  6.54 0.00  $\pm$  0.00 0.00  $\pm$  0.00 34.86  $\pm$  10.12 24.65  $\pm$  4.02 15.52  $\pm$  5.30 22.74  $\pm$  5.03
REPLM (random all) 0.00  $\pm$  0.00 16.99  $\pm$  3.11 0.00  $\pm$  0.00 0.00  $\pm$  0.00 33.58  $\pm$  3.41 25.84  $\pm$  5.77 14.65  $\pm$  4.25 25.44  $\pm$  4.98
REPLM (best $\ominus$) 31.58 45.54 0.00 0.00 52.38 42.35 20.00 20.11
REPLM (best $\oplus$) 31.58 43.40 0.00 0.00 68.29 34.48 29.71 16.48
REPLM 26.09 35.80 0.00 0.00 55.17 28.81 30.15 28.22
REPLM (params adj) 26.09 35.58 0.00 0.00 60.87 28.81 30.61 29.27
Appendix I Full Results on Comparison Against External Knowledge
Tables 34 to 41 provide the detailed results of relation extraction of all methods on all relation types. The evaluation is done based on checking the correctness of extracted relations against both the human annotations of the dev documents and an external knowledge base.
Overall, we see that our our REPLM and REPLM (params adj) outperforms REBEL in most relations, when further including the relations from an external knowledge base. The performance improvement becomes more striking for the relation types having a large number of knowledge triplets. For instance, for the relation P17 (country), our REPLM achieves an F1 score of 56.14, whereas REBEL-sent can achieve less than the half of the performance, F1 score of 21.17. We want to highlight that, these methods performed at similar level when compared against only human-annotations as ground-truth (REPLM with 27.74 F1 score vs. REBEL-sent with 25.73 F1 score, see Sec. H). Therefore, it shows the importance of evaluating against an external knowledge base. Even a larger performance gap can be found in relation P27 (country of citizenship), our REPLM and REPLM (params adj) achieve the F1 scores 39.52 and 46.15, respectively, whereas REBEL and REBEL-sent can achieve only F1 scores 3.47 and 8.24. These results further confirm the effectiveness of our REPLM framework, agreeing with the results from the earlier sections.
Table 34: Full results in comparison against external knowledge (Section 6.2). Shown are F1 scores on each relation. (Part 1 / 8)
Method P6 P17 P19 P20 P22 P25 P26 P27 P30 P31 P35 P36
REBEL 0.00 17.08 37.46 13.92 35.09 30.77 35.86 3.47 3.22 4.80 0.00 15.34
REBEL-sent 2.13 21.17 34.55 14.29 27.74 26.32 30.66 7.72 8.24 2.81 2.78 16.90
REPLM 14.81 56.14 33.83 10.90 3.86 9.16 6.40 39.52 49.16 9.78 5.85 14.63
REPLM (params adj) 14.04 55.32 53.67 37.84 12.68 10.31 4.08 46.15 39.22 13.89 13.86 26.94
Table 35: Full results in comparison against external knowledge (Section 6.2). Shown are F1 scores on each relation. (Part 2 / 8)
Method P37 P39 P40 P50 P54 P57 P58 P69 P86 P102 P108 P112
REBEL 0.00 30.11 34.97 43.27 37.84 48.05 16.90 35.79 31.25 47.13 27.91 15.38
REBEL-sent 0.00 13.39 28.07 40.89 24.00 41.56 2.82 5.23 15.38 57.58 20.00 10.31
REPLM 8.79 2.50 2.75 14.15 45.04 38.82 24.78 37.76 19.48 37.04 6.01 3.12
REPLM (params adj) 18.60 15.00 4.97 30.84 45.04 38.32 25.23 35.23 25.64 37.78 22.22 12.96
Table 36: Full results in comparison against external knowledge (Section 6.2). Shown are F1 scores on each relation. (Part 3 / 8)
Method P118 P123 P127 P131 P136 P137 P140 P150 P155 P156 P159 P161
REBEL 29.09 26.67 12.28 28.54 4.29 12.50 4.12 29.01 6.67 8.06 13.86 42.52
REBEL-sent 36.99 23.44 12.31 26.03 2.37 5.06 12.70 33.80 7.91 12.31 21.67 21.20
REPLM 34.55 25.24 6.61 38.54 19.40 9.09 0.00 43.21 9.52 4.88 21.88 24.62
REPLM (params adj) 33.63 26.92 11.03 42.39 13.33 9.52 8.55 47.81 10.98 11.70 23.85 32.56
Table 37: Full results in comparison against external knowledge (Section 6.2). Shown are F1 scores on each relation. (Part 4 / 8)
Method P162 P166 P170 P171 P172 P175 P176 P178 P179 P190 P194 P205
REBEL 10.20 31.37 25.00 10.81 0.00 49.91 15.62 30.43 0.00 0.00 9.09 5.56
REBEL-sent 6.38 36.21 24.44 6.45 0.00 46.49 32.18 32.06 0.00 0.00 14.14 15.58
REPLM 11.04 5.94 6.90 27.91 12.28 42.04 24.76 1.06 0.62 6.85 9.30 12.50
REPLM (params adj) 10.48 26.79 10.34 26.32 21.97 42.16 24.72 16.36 19.51 6.18 14.43 14.46
Table 38: Full results in comparison against external knowledge (Section 6.2). Shown are F1 scores on each relation. (Part 5 / 8)
Method P206 P241 P264 P272 P276 P279 P355 P361 P364 P400 P403 P449
REBEL 28.88 27.59 44.39 38.71 18.03 5.00 10.67 27.42 0.00 37.38 36.11 0.00
REBEL-sent 24.91 12.90 27.37 40.58 18.00 2.28 14.37 12.41 0.00 34.00 36.89 0.00
REPLM 4.41 3.96 0.00 4.61 9.01 19.72 6.90 16.35 16.25 32.18 23.73 48.65
REPLM (params adj) 4.32 58.97 29.06 27.35 9.35 15.24 9.52 13.30 37.91 32.18 23.08 44.44
Table 39: Full results in comparison against external knowledge (Section 6.2). Shown are F1 scores on each relation. (Part 6 / 8)
Method P463 P488 P495 P527 P551 P569 P570 P571 P576 P577 P580 P582
REBEL 32.03 15.38 3.47 29.70 7.14 50.08 38.79 44.90 0.00 34.01 19.51 0.00
REBEL-sent 24.58 5.88 4.50 16.55 10.81 50.82 43.56 27.59 0.00 39.40 14.71 0.00
REPLM 9.33 0.00 12.52 8.43 5.56 57.76 38.81 29.67 0.97 21.96 6.90 0.86
REPLM (params adj) 15.92 16.95 19.80 20.00 5.00 62.15 46.31 31.20 4.00 42.14 7.48 5.61
Table 40: Full results in comparison against external knowledge (Section 6.2). Shown are F1 scores on each relation. (Part 7 / 8)
Method P585 P607 P674 P676 P706 P710 P737 P740 P749 P800 P807 P840
REBEL 5.50 38.78 25.64 0.00 11.43 16.53 0.00 11.76 5.26 29.31 0.00 0.00
REBEL-sent 3.35 34.36 19.86 0.00 8.62 13.79 0.00 12.90 14.77 24.71 0.00 0.00
REPLM 21.43 26.14 18.02 3.62 7.62 2.55 0.00 25.93 0.00 8.89 2.98 3.95
REPLM (params adj) 20.00 24.86 15.24 14.71 7.62 26.67 6.67 27.40 6.32 17.35 15.24 4.94
Table 41: Full results in comparison against external knowledge (Section 6.2). Shown are F1 scores on each relation. (Part 8 / 8)
Method P937 P1001 P1056 P1198 P1336 P1344 P1365 P1366 P1376 P1412 P1441 P3373
REBEL 0.00 11.03 0.00 0.00 0.00 0.00 6.25 3.70 7.06 0.00 28.07 39.58
REBEL-sent 7.06 3.37 0.00 0.00 0.00 0.00 10.81 4.12 4.61 0.00 19.29 32.93
REPLM 4.42 12.56 0.00 0.00 0.00 49.52 0.00 0.00 36.36 34.22 15.71 7.41
REPLM (params adj) 12.35 17.83 0.00 1.19 13.95 47.27 0.00 0.00 37.85 12.03 15.83 8.78
Appendix J Sensitivity Analysis of our REPLM Framework
In our original REPLM framework, we used a fixed temperature $\tau$ and a fixed probability threshold $\theta$ as our REPLM does not require a human-annotated training documents to tune the hyperparameters. Tables 42 to 47 show the sensitivity of REPLM on various hyperparameter selections for all relation types individually. Overall, we observe that our REPLM framework is robust to different hyperparameter: the performance remains at the same level for most variations. (Only exception is $\theta=0.5$, where such high probability threshold results in discarding the correctly extracted relations. This leads to much lower recall, and therefore, F1 score.) This also explains why the performance gap between REPLM and REPLM (params adj) is small.
Table 42: Sensitivity of our REPLM framework predictions w.r.t $\tau$ and $\theta$. Shown are F1 scores on each relation. (Part 1 / 6)
$\tau$ $\theta$ P6 P17 P19 P20 P22 P25 P26 P27 P30 P31 P35 P36 P37 P39 P40 P50
0.01 0.01 33.11 25.88 57.28 41.05 16.11 22.58 35.96 29.08 36.69 11.51 30.88 45.95 22.22 8.70 17.87 40.00
0.01 0.05 34.07 26.13 60.31 43.43 17.30 20.83 36.89 30.39 38.43 12.61 29.03 51.61 27.37 9.52 16.22 42.05
0.01 0.10 36.67 25.36 64.09 47.50 12.82 14.63 36.27 31.77 32.69 12.77 32.73 53.33 30.95 0.00 16.67 38.67
0.01 0.20 40.40 24.55 71.03 54.96 12.60 20.00 38.32 32.39 31.46 14.63 34.69 48.00 27.40 0.00 16.18 37.50
0.01 0.50 24.62 17.25 76.26 57.78 14.12 10.00 35.48 25.56 26.49 12.70 30.14 42.86 10.91 0.00 17.09 33.57
0.05 0.01 31.06 25.78 55.94 40.20 17.65 22.54 36.89 28.29 33.87 10.39 29.17 45.00 21.85 7.14 17.09 41.41
0.05 0.05 31.25 26.86 56.34 40.20 17.94 22.95 36.51 28.81 38.20 11.94 29.37 46.58 25.00 8.33 17.65 40.00
0.05 0.10 35.56 27.63 60.10 41.05 15.54 28.00 38.01 30.75 39.46 14.16 33.33 52.46 27.91 9.09 17.83 41.18
0.05 0.20 32.65 26.12 69.91 52.24 20.80 16.67 40.24 34.45 34.09 15.58 36.00 52.83 29.73 12.50 17.75 41.98
0.05 0.50 16.95 13.68 77.94 57.47 13.33 11.11 31.78 24.81 22.07 7.02 31.43 46.15 11.32 0.00 12.24 33.59
0.10 0.01 31.06 25.75 55.94 40.20 17.65 22.54 36.89 28.29 33.65 10.32 29.17 45.00 21.85 7.14 17.03 41.41
0.10 0.05 31.25 26.85 56.34 40.20 17.47 22.22 37.04 28.68 36.36 11.76 29.58 45.95 25.24 8.00 17.59 41.78
0.10 0.10 32.99 26.16 70.34 53.03 19.83 16.67 40.25 34.65 34.48 15.79 35.05 52.83 29.73 12.50 18.63 41.25
0.10 0.20 34.78 27.74 59.06 40.41 17.17 28.00 37.17 30.71 39.11 14.29 32.56 52.46 27.91 8.70 17.56 41.12
0.10 0.50 16.95 13.63 77.49 58.14 13.33 11.11 32.08 24.84 22.07 7.02 30.99 46.15 11.32 0.00 10.31 33.59
0.20 0.01 31.06 25.72 55.94 40.20 17.65 22.54 36.89 28.29 33.65 10.32 29.17 45.00 21.85 7.14 17.03 41.41
0.20 0.05 31.25 26.78 56.21 40.20 18.26 22.22 37.19 28.66 34.91 11.76 29.58 45.33 25.24 8.00 17.53 41.59
0.20 0.10 34.53 27.68 59.20 40.41 18.00 28.00 37.00 30.56 38.94 14.29 31.82 51.61 28.24 9.09 17.16 43.52
0.20 0.20 33.33 26.39 70.99 53.12 18.64 11.43 41.56 34.56 34.48 16.00 35.42 52.83 30.14 13.33 20.13 41.51
0.20 0.50 20.00 13.71 77.49 58.14 13.33 11.11 33.64 24.86 20.83 7.02 30.99 46.15 11.32 0.00 10.20 33.59
0.50 0.01 31.06 25.73 55.94 40.20 17.65 22.54 36.89 28.28 33.65 10.32 29.17 45.00 21.85 7.14 17.03 41.41
0.50 0.05 31.25 26.73 56.21 40.20 18.26 22.58 37.19 28.65 34.78 11.76 29.58 45.95 25.49 8.00 17.53 41.59
0.50 0.10 34.04 27.77 59.35 40.21 17.91 28.00 37.00 30.54 38.94 14.04 31.58 51.61 27.91 9.09 17.23 43.12
0.50 0.20 29.79 26.23 70.99 53.97 18.49 11.43 39.74 34.55 33.33 15.79 35.42 52.83 30.14 13.33 20.25 41.77
0.50 0.50 20.00 13.96 77.49 59.77 13.33 11.11 33.64 24.98 20.83 7.02 23.19 46.15 11.32 0.00 10.20 33.59
1.00 0.01 31.06 25.73 55.94 40.20 17.65 22.54 36.89 28.28 33.65 10.32 29.17 45.00 21.85 7.14 17.03 41.41
1.00 0.05 31.25 26.73 56.21 40.20 18.26 22.22 37.19 28.65 34.91 11.76 29.58 45.95 25.24 8.00 17.53 41.59
1.00 0.10 33.09 27.85 59.35 40.21 18.72 28.00 37.00 30.54 38.77 14.04 31.34 51.61 27.91 9.09 17.10 43.12
1.00 0.20 29.79 26.18 71.21 53.97 18.80 11.43 39.74 34.64 33.33 15.79 35.42 52.83 30.56 13.33 20.38 41.77
1.00 0.50 20.00 13.97 77.49 59.77 13.33 11.11 32.08 24.81 20.83 7.02 23.19 46.15 11.32 0.00 10.20 33.59
Table 43: Sensitivity of our REPLM framework predictions w.r.t $\tau$ and $\theta$. Shown are F1 scores on each relation. (Part 2 / 6)
$\tau$ $\theta$ P54 P57 P58 P69 P86 P102 P108 P112 P118 P123 P127 P131 P136 P137 P140 P150
0.01 0.01 44.71 42.55 30.00 56.39 52.54 39.43 29.73 30.14 35.16 32.47 19.28 29.07 21.05 15.38 16.17 32.25
0.01 0.05 45.26 43.17 31.58 60.29 50.45 38.71 31.34 28.57 39.47 32.17 20.13 28.64 25.81 17.95 15.03 32.92
0.01 0.10 42.81 43.48 33.33 58.59 49.06 42.92 31.15 33.33 40.88 31.34 20.00 27.52 20.69 13.51 13.70 34.45
0.01 0.20 39.86 45.41 35.82 59.67 39.58 45.60 31.48 35.56 42.28 29.75 19.67 24.54 17.39 16.13 13.79 33.10
0.01 0.50 41.53 47.20 37.93 61.44 19.18 41.77 27.78 36.84 37.65 30.43 18.56 15.82 11.11 8.16 6.38 24.12
0.05 0.01 43.06 42.28 29.63 56.77 52.54 39.04 30.07 26.83 34.97 32.34 19.69 28.97 24.39 14.00 15.44 32.13
0.05 0.05 44.06 42.98 29.63 56.77 52.54 39.72 30.46 29.33 35.96 31.90 20.11 29.35 27.03 14.58 15.02 32.86
0.05 0.10 45.45 44.44 31.58 55.86 51.79 41.38 31.65 31.25 40.25 32.65 22.08 30.02 18.75 17.95 14.55 34.27
0.05 0.20 43.42 47.42 35.82 64.41 38.71 46.07 34.00 40.91 45.90 35.20 20.87 26.60 16.00 17.54 13.45 33.77
0.05 0.50 35.71 48.05 38.60 60.00 17.14 35.71 25.71 24.24 37.04 30.23 17.78 13.51 12.50 4.35 4.55 20.25
0.10 0.01 43.06 42.28 29.63 56.77 52.54 39.04 30.07 26.51 34.97 32.34 19.59 28.95 24.39 14.00 15.50 32.11
0.10 0.05 43.68 42.98 29.63 56.77 52.54 39.86 30.46 28.57 35.96 32.73 20.65 29.39 27.03 14.43 14.88 32.72
0.10 0.10 43.42 47.12 36.36 64.74 38.71 45.90 35.42 37.21 46.28 35.77 22.41 26.68 16.67 17.24 13.68 33.74
0.10 0.20 45.78 43.70 30.77 56.25 51.33 42.26 31.88 31.43 39.75 32.47 21.52 29.89 18.18 17.72 14.29 34.36
0.10 0.50 36.28 47.06 39.29 59.06 17.14 36.23 25.71 24.24 37.04 30.59 19.78 13.55 12.50 4.35 4.60 20.08
0.20 0.01 43.06 42.28 29.63 56.77 52.54 39.04 30.07 26.51 34.97 32.34 19.59 28.95 24.39 14.00 15.44 32.11
0.20 0.05 43.68 42.80 29.63 56.77 52.54 39.86 30.46 28.57 35.96 32.53 20.54 29.27 27.03 14.58 14.75 32.77
0.20 0.10 45.78 43.51 30.38 56.25 49.12 42.38 32.17 31.43 39.75 33.99 21.79 30.07 18.75 17.72 13.25 34.15
0.20 0.20 43.42 45.74 36.92 65.12 35.16 43.43 36.56 37.21 47.06 37.61 22.61 26.65 17.39 17.24 13.68 34.10
0.20 0.50 35.96 47.06 40.74 58.11 17.14 36.50 25.71 24.24 39.02 30.59 20.00 13.31 12.50 4.44 4.60 19.58
0.50 0.01 43.06 42.28 29.63 56.77 52.54 39.04 30.07 26.51 34.97 32.34 19.59 28.95 24.39 14.00 15.44 32.11
0.50 0.05 43.68 42.62 29.63 56.77 52.54 39.58 30.46 28.57 35.96 32.53 20.54 29.29 27.03 14.58 14.75 32.73
0.50 0.10 45.65 43.15 30.38 56.50 47.79 42.07 32.17 31.43 39.51 33.99 21.79 29.89 24.24 17.50 13.10 33.90
0.50 0.20 44.04 45.99 36.92 65.12 35.56 43.43 34.78 37.21 47.46 36.84 23.01 26.47 16.67 17.54 13.56 34.02
0.50 0.50 35.96 48.05 40.74 58.11 17.14 36.50 25.71 29.41 39.02 30.59 20.00 13.32 12.50 4.44 4.60 19.89
1.00 0.01 43.06 42.28 29.63 56.77 52.54 39.04 30.07 26.19 34.97 32.34 19.59 28.95 24.39 14.00 15.44 32.11
1.00 0.05 43.68 42.62 29.63 56.77 52.54 39.58 30.46 28.57 35.96 32.53 20.54 29.30 27.03 14.58 14.61 32.71
1.00 0.10 45.51 43.15 30.38 57.14 47.79 42.07 32.17 31.43 39.26 33.99 21.79 29.79 24.24 17.50 13.10 33.92
1.00 0.20 44.20 45.99 36.92 65.50 35.56 42.53 34.78 37.21 47.46 36.84 23.21 26.37 16.67 17.54 13.56 33.75
1.00 0.50 35.96 48.05 40.74 58.11 17.14 35.29 25.71 29.41 39.02 30.59 20.00 13.32 12.50 8.70 4.60 19.89
Table 44: Sensitivity of our REPLM framework predictions w.r.t $\tau$ and $\theta$. Shown are F1 scores on each relation. (Part 3 / 6)
$\tau$ $\theta$ P155 P156 P159 P161 P162 P166 P170 P171 P172 P175 P176 P178 P179 P190 P194 P205
0.01 0.01 15.84 26.44 41.58 45.04 24.06 27.50 16.81 14.63 31.37 50.90 27.03 32.51 25.00 66.67 26.21 23.38
0.01 0.05 15.57 28.17 44.09 46.60 24.59 27.40 17.48 10.81 35.44 51.75 28.57 32.46 27.64 66.67 27.87 24.62
0.01 0.10 17.39 26.23 43.68 45.96 23.64 27.48 14.74 11.43 37.68 52.24 28.99 31.82 26.55 66.67 25.69 27.12
0.01 0.20 13.79 27.08 43.87 43.17 16.28 30.36 13.33 12.12 30.51 50.17 25.00 32.26 26.42 66.67 24.74 24.49
0.01 0.50 14.12 24.24 35.59 35.53 8.96 30.77 11.76 14.81 26.09 43.15 9.09 19.64 23.26 66.67 17.39 5.13
0.05 0.01 14.29 25.26 39.09 43.73 24.64 26.51 17.14 14.63 28.32 49.94 28.57 31.78 24.83 66.67 24.20 23.53
0.05 0.05 14.29 26.82 40.76 44.18 24.82 27.16 18.18 14.63 31.58 50.86 28.95 32.23 25.00 66.67 25.50 25.32
0.05 0.10 16.47 31.21 46.24 46.21 24.00 27.27 14.55 14.63 32.91 51.99 28.17 31.02 26.98 66.67 29.27 23.88
0.05 0.20 11.97 27.66 44.30 47.86 26.37 34.23 17.72 11.43 30.77 53.27 34.48 32.89 22.43 66.67 21.51 16.00
0.05 0.50 14.81 22.58 35.09 36.80 9.38 27.59 8.16 15.38 15.00 39.39 4.76 18.37 20.25 66.67 12.70 5.26
0.10 0.01 14.29 25.26 39.09 43.73 24.64 26.51 17.14 14.63 28.32 49.94 28.57 31.78 24.83 66.67 24.20 23.53
0.10 0.05 14.08 26.52 40.57 44.03 24.82 27.16 17.52 14.63 30.30 50.55 28.95 31.92 24.83 66.67 25.50 25.00
0.10 0.10 12.28 27.37 42.31 47.95 27.59 34.55 17.50 11.43 29.63 53.00 31.58 32.43 23.30 66.67 23.91 16.33
0.10 0.20 16.09 29.93 45.03 45.23 23.81 26.42 15.52 14.63 32.91 52.19 28.17 31.75 27.27 66.67 30.65 25.35
0.10 0.50 14.81 19.67 37.93 36.90 9.38 27.59 12.50 16.00 10.26 38.94 9.30 19.57 20.00 66.67 15.62 5.13
0.20 0.01 14.29 25.26 39.09 43.73 24.64 26.35 17.14 14.63 28.32 49.94 28.57 31.78 24.83 66.67 24.20 23.53
0.20 0.05 13.95 26.52 40.57 44.03 24.82 27.16 17.52 14.63 30.30 50.61 28.95 31.92 24.83 66.67 25.50 25.00
0.20 0.10 15.56 30.07 44.33 45.23 23.62 26.42 15.00 14.63 32.50 51.91 30.14 32.46 27.07 66.67 30.40 25.35
0.20 0.20 12.28 28.57 41.83 47.71 30.23 33.64 17.07 11.76 26.92 52.88 32.14 30.99 22.00 66.67 21.98 16.00
0.20 0.50 14.63 20.00 36.52 35.82 9.38 27.59 12.50 16.00 10.26 38.22 9.30 19.35 20.00 66.67 15.62 5.13
0.50 0.01 14.29 25.26 39.09 43.73 24.64 26.35 17.14 14.63 28.32 49.94 28.57 31.78 24.83 66.67 24.20 23.53
0.50 0.05 13.95 26.52 40.76 44.03 24.82 26.99 17.52 14.63 30.30 50.49 28.95 31.78 24.83 66.67 25.68 24.69
0.50 0.10 15.38 30.07 43.88 45.23 23.81 27.50 14.75 14.63 32.10 51.84 29.73 32.12 27.27 66.67 30.16 25.35
0.50 0.20 12.28 28.89 42.11 48.37 30.23 32.08 17.50 11.76 26.92 52.76 32.14 31.88 20.62 66.67 21.98 12.24
0.50 0.50 14.63 20.00 36.52 35.71 9.38 27.59 13.04 16.00 10.26 37.58 9.30 19.35 20.00 66.67 15.62 5.13
1.00 0.01 14.29 25.26 39.09 43.73 24.64 26.35 17.14 14.63 28.32 49.94 28.57 31.78 24.83 66.67 24.20 23.53
1.00 0.05 13.95 26.52 40.76 44.10 24.82 26.99 17.52 14.63 30.30 50.49 28.95 31.78 24.83 66.67 25.68 24.69
1.00 0.10 15.47 30.97 44.10 45.23 23.81 27.50 14.52 14.63 32.50 51.78 29.73 31.96 27.27 66.67 30.16 25.00
1.00 0.20 12.28 28.89 41.83 48.26 30.23 32.08 17.50 11.76 26.92 52.94 32.14 31.88 20.83 66.67 21.98 12.24
1.00 0.50 14.81 20.00 36.52 36.20 9.38 31.11 12.77 16.00 10.26 37.58 9.30 19.35 20.00 66.67 15.62 5.13
Table 45: Sensitivity of our REPLM framework predictions w.r.t $\tau$ and $\theta$. Shown are F1 scores on each relation. (Part 4 / 6)
$\tau$ $\theta$ P206 P241 P264 P272 P276 P279 P355 P361 P364 P400 P403 P449 P463 P488 P495 P527
0.01 0.01 19.78 41.90 43.53 37.04 26.60 11.11 24.66 22.50 34.29 32.26 28.57 33.71 38.26 14.29 21.37 20.21
0.01 0.05 19.74 42.86 43.43 35.62 27.16 8.57 28.07 25.66 36.36 33.33 28.95 34.88 39.68 16.33 22.54 21.41
0.01 0.10 16.78 45.45 43.21 37.14 29.58 6.90 26.92 26.53 28.07 32.00 23.53 31.71 38.94 13.95 21.82 21.97
0.01 0.20 11.57 48.19 39.71 37.29 29.03 8.33 26.67 26.38 23.53 35.09 26.23 33.77 38.38 16.22 23.73 20.88
0.01 0.50 4.12 47.62 34.12 34.62 22.45 9.52 27.03 21.05 22.22 27.08 20.41 33.33 34.62 19.35 18.57 21.19
0.05 0.01 19.79 44.04 44.10 37.65 23.58 12.61 20.00 21.53 33.80 32.05 26.97 35.16 35.29 12.90 21.68 17.44
0.05 0.05 21.11 42.99 44.33 36.59 25.71 13.48 22.22 22.80 34.29 32.26 27.27 35.56 37.62 13.33 21.76 19.68
0.05 0.10 20.38 45.36 45.10 36.36 30.12 9.38 26.67 25.23 38.10 35.21 27.40 36.14 43.02 14.81 21.75 22.17
0.05 0.20 17.24 52.63 44.66 40.68 35.48 8.51 23.81 30.72 23.53 34.19 27.12 36.11 43.52 20.00 23.76 23.49
0.05 0.50 6.45 53.12 33.44 36.73 20.45 0.00 18.18 20.54 22.22 29.21 21.28 33.33 31.72 15.38 22.99 18.96
0.10 0.01 19.79 44.04 44.10 37.65 23.58 12.61 20.00 21.43 33.80 32.05 26.97 35.16 35.29 12.90 21.68 17.73
0.10 0.05 20.77 44.04 44.41 36.14 25.00 13.64 21.43 22.50 34.29 32.26 27.27 35.16 36.77 13.11 21.96 19.68
0.10 0.10 17.09 52.63 45.93 42.11 32.79 8.89 23.26 30.19 24.00 35.40 28.07 39.39 42.71 20.00 24.30 24.66
0.10 0.20 20.51 46.00 45.54 36.36 30.12 9.09 22.22 25.39 39.34 34.01 27.03 38.10 42.07 14.81 22.41 21.74
0.10 0.50 4.35 53.97 33.54 37.50 20.69 0.00 18.18 20.54 22.22 29.55 21.28 33.33 32.39 8.33 22.22 18.54
0.20 0.01 19.79 44.04 44.10 37.65 23.58 12.61 20.00 21.39 33.80 32.05 26.97 35.16 35.29 12.90 21.68 17.73
0.20 0.05 20.88 44.04 44.41 35.71 25.12 13.64 21.43 21.98 34.78 32.26 27.27 35.16 36.77 12.90 22.06 19.56
0.20 0.10 19.74 45.54 45.54 35.90 29.94 8.82 22.22 25.11 39.34 33.56 29.33 37.65 41.18 15.09 22.67 21.85
0.20 0.20 17.24 52.63 46.75 42.11 32.52 0.00 23.81 29.30 24.00 37.04 28.07 37.50 41.88 15.79 24.68 24.74
0.20 0.50 4.35 53.97 32.81 37.50 20.69 0.00 18.18 20.54 22.22 25.58 21.28 33.33 32.39 8.33 22.05 18.45
0.50 0.01 19.79 44.04 44.10 37.65 23.58 12.73 20.00 21.43 33.80 32.05 26.97 35.16 35.29 12.90 21.68 17.73
0.50 0.05 20.88 44.04 44.41 35.71 25.12 13.64 21.43 21.66 34.78 32.26 27.27 35.16 36.77 12.90 22.06 19.60
0.50 0.10 19.74 45.54 45.37 35.90 30.49 8.96 21.54 25.17 39.34 33.33 28.95 38.10 41.33 15.09 22.67 21.80
0.50 0.20 17.24 52.63 46.63 42.11 32.52 0.00 23.81 29.39 24.49 37.38 28.07 38.10 41.05 15.79 24.68 24.74
0.50 0.50 4.35 53.97 31.85 40.82 20.69 0.00 18.18 21.33 22.22 23.53 21.28 33.33 32.39 7.69 22.05 18.54
1.00 0.01 19.79 44.04 44.10 37.65 23.58 12.73 20.00 21.43 33.80 32.05 26.97 35.16 35.29 12.90 21.68 17.73
1.00 0.05 20.88 44.04 44.41 35.71 25.12 13.79 21.43 21.66 34.78 32.26 27.27 35.16 36.77 12.90 22.06 19.64
1.00 0.10 20.00 45.54 45.45 35.90 30.67 8.96 21.21 25.33 39.34 32.89 28.95 38.10 41.48 14.81 22.63 21.70
1.00 0.20 17.24 52.63 46.75 42.11 32.26 0.00 23.81 28.85 24.49 37.74 28.07 38.10 42.71 15.79 24.81 24.91
1.00 0.50 4.35 53.97 31.85 40.82 20.69 0.00 18.18 21.33 22.22 23.53 21.28 33.33 32.39 7.69 22.05 18.54
Table 46: Sensitivity of our REPLM framework predictions w.r.t $\tau$ and $\theta$. Shown are F1 scores on each relation. (Part 5 / 6)
$\tau$ $\theta$ P551 P569 P570 P571 P576 P577 P580 P582 P585 P607 P674 P676 P706 P710 P737 P740
0.01 0.01 26.67 55.59 42.86 37.04 14.55 49.90 23.33 38.10 25.93 31.23 30.08 47.06 21.77 26.02 0.00 33.33
0.01 0.05 30.77 58.03 44.24 40.31 15.05 52.24 26.42 40.74 25.81 31.93 24.79 57.14 24.81 29.09 0.00 30.77
0.01 0.10 36.36 59.68 45.71 44.13 15.38 52.85 27.59 42.55 28.92 31.63 22.61 57.14 22.81 29.70 0.00 28.57
0.01 0.20 36.36 61.61 49.16 46.25 9.84 51.57 32.84 43.90 28.57 31.31 24.53 46.15 16.67 27.37 0.00 31.25
0.01 0.50 22.22 62.01 50.36 43.98 12.00 46.79 35.56 36.36 22.22 25.00 23.66 46.15 13.51 24.69 0.00 41.67
0.05 0.01 25.00 54.43 41.94 34.61 12.60 47.98 21.37 36.36 27.03 29.89 30.00 47.06 19.51 24.43 0.00 31.11
0.05 0.05 25.00 54.78 42.48 36.99 14.16 50.35 23.53 38.71 29.13 30.77 30.66 47.06 20.65 25.81 0.00 34.15
0.05 0.10 28.57 56.29 44.87 40.89 17.98 52.46 28.28 46.15 28.26 33.87 30.40 50.00 24.43 30.19 0.00 32.43
0.05 0.20 33.33 61.47 49.79 45.66 19.05 54.40 34.92 43.90 27.78 27.69 23.64 46.15 17.39 23.66 0.00 32.26
0.05 0.50 28.57 61.18 50.36 44.55 8.33 46.69 28.57 26.67 26.92 23.45 18.18 46.15 14.08 18.42 0.00 38.10
0.10 0.01 25.00 54.43 41.94 34.61 12.60 47.94 21.37 36.36 27.03 29.89 30.00 47.06 19.51 24.06 0.00 31.11
0.10 0.05 26.67 54.71 42.41 36.75 13.56 50.05 23.33 37.50 29.13 30.22 30.22 47.06 20.25 25.60 0.00 34.15
0.10 0.10 33.33 62.10 50.21 46.15 19.35 53.90 35.48 45.00 26.47 26.04 23.85 46.15 17.78 23.66 0.00 33.33
0.10 0.20 28.57 55.73 44.56 39.71 17.78 52.39 29.17 43.14 30.11 33.20 31.25 47.06 23.02 30.19 0.00 31.58
0.10 0.50 28.57 61.29 50.36 43.44 8.51 46.96 28.57 26.67 27.45 23.29 17.78 46.15 14.08 18.67 0.00 40.00
0.20 0.01 25.00 54.43 41.94 34.61 12.60 47.98 21.37 36.36 27.03 29.89 30.00 47.06 19.51 24.06 0.00 31.11
0.20 0.05 26.67 54.64 42.35 36.83 13.68 49.85 23.14 36.92 29.13 30.22 30.22 47.06 20.25 25.60 0.00 34.15
0.20 0.10 26.67 55.51 44.48 39.42 17.58 52.61 28.87 44.00 30.11 32.94 31.01 47.06 22.86 30.19 0.00 32.43
0.20 0.20 33.33 61.77 50.54 45.95 19.35 53.42 35.48 45.00 26.47 26.18 23.30 46.15 17.78 23.66 0.00 33.33
0.20 0.50 28.57 61.18 50.00 43.44 8.51 47.02 28.57 25.81 27.45 21.92 17.78 46.15 14.08 18.67 0.00 40.00
0.50 0.01 25.00 54.43 41.94 34.61 12.60 47.94 21.37 36.36 27.03 29.89 30.00 47.06 19.51 24.06 0.00 31.11
0.50 0.05 26.67 54.64 42.35 36.91 13.56 49.75 22.95 36.92 29.13 30.22 30.43 47.06 20.13 25.60 0.00 34.15
0.50 0.10 26.67 55.44 44.41 38.94 17.39 52.38 26.53 43.14 30.11 32.81 31.01 47.06 23.02 30.19 0.00 32.43
0.50 0.20 33.33 61.11 50.66 46.10 19.05 53.37 35.48 45.00 26.47 25.26 23.30 46.15 17.98 23.91 0.00 33.33
0.50 0.50 28.57 61.29 50.85 44.14 8.70 46.94 28.57 25.81 26.92 21.77 17.78 46.15 14.08 18.67 0.00 40.00
1.00 0.01 25.00 54.43 41.94 34.61 12.60 47.94 21.37 36.36 27.03 29.89 30.00 47.06 19.51 24.06 0.00 31.11
1.00 0.05 26.67 54.57 42.35 36.83 13.56 49.75 22.95 36.92 29.13 30.22 30.43 47.06 20.00 25.60 0.00 34.15
1.00 0.10 26.67 55.44 44.33 38.94 17.58 52.22 26.53 43.14 30.11 32.81 30.77 47.06 23.02 30.19 0.00 32.43
1.00 0.20 33.33 61.21 50.66 46.78 19.05 53.19 35.48 41.03 26.47 25.26 23.30 46.15 17.98 23.91 0.00 33.33
1.00 0.50 28.57 61.29 50.85 44.14 8.70 46.94 28.57 25.81 26.92 21.77 17.78 46.15 14.08 18.67 0.00 40.00
Table 47: Sensitivity of our REPLM framework predictions w.r.t $\tau$ and $\theta$. Shown are F1 scores on each relation. (Part 6 / 6)
$\tau$ $\theta$ P749 P800 P807 P840 P937 P1001 P1056 P1198 P1336 P1344 P1365 P1366 P1376 P1412 P1441 P3373
0.01 0.01 22.02 27.94 50.00 43.24 27.27 22.75 0.00 66.67 23.08 34.15 0.00 6.25 56.67 28.35 30.41 29.27
0.01 0.05 20.22 28.80 50.00 41.18 30.99 24.42 0.00 100.00 25.00 36.36 0.00 8.33 57.14 28.83 27.41 28.96
0.01 0.10 21.05 28.57 66.67 40.00 33.33 25.85 0.00 100.00 27.27 34.59 0.00 9.09 52.00 28.00 27.12 25.96
0.01 0.20 24.62 25.81 66.67 48.00 35.29 29.03 0.00 100.00 21.05 37.04 0.00 13.33 54.55 31.58 27.71 24.74
0.01 0.50 20.00 26.67 0.00 30.00 27.78 23.08 0.00 0.00 14.29 41.38 0.00 0.00 60.61 32.79 17.81 16.57
0.05 0.01 21.43 27.74 50.00 46.15 26.67 21.10 0.00 66.67 19.35 34.12 0.00 4.35 53.97 26.09 31.49 29.60
0.05 0.05 20.97 28.15 50.00 42.11 25.29 23.30 0.00 100.00 23.08 34.12 0.00 5.88 54.84 27.07 29.36 29.60
0.05 0.10 19.35 27.87 50.00 36.36 29.33 28.05 0.00 100.00 26.09 36.02 0.00 10.00 56.14 30.09 30.93 27.85
0.05 0.20 25.81 30.77 66.67 30.77 26.92 30.65 0.00 100.00 12.50 37.50 0.00 0.00 60.87 32.50 24.54 26.26
0.05 0.50 17.39 27.40 0.00 21.05 21.43 16.84 0.00 0.00 15.38 41.86 0.00 0.00 58.06 24.14 18.84 19.63
0.10 0.01 21.43 27.74 50.00 46.15 26.67 21.10 0.00 66.67 19.35 34.12 0.00 4.35 53.97 26.09 31.49 29.60
0.10 0.05 21.37 27.94 50.00 42.11 25.29 23.81 0.00 100.00 21.43 34.12 0.00 5.26 54.84 26.87 30.00 29.60
0.10 0.10 22.58 31.82 66.67 30.77 21.74 30.40 0.00 100.00 12.50 37.84 0.00 0.00 62.22 33.33 23.75 25.13
0.10 0.20 24.49 29.27 50.00 34.29 28.57 28.57 0.00 100.00 26.09 35.80 0.00 0.00 55.17 28.81 30.15 28.22
0.10 0.50 17.39 27.03 0.00 21.05 20.69 16.84 0.00 0.00 15.38 41.38 0.00 0.00 56.25 24.56 17.65 19.75
0.20 0.01 21.43 27.74 50.00 46.15 26.67 21.10 0.00 66.67 19.35 34.12 0.00 4.35 53.97 26.09 31.49 29.60
0.20 0.05 22.73 27.94 50.00 42.11 25.29 23.81 0.00 100.00 21.43 34.12 0.00 5.00 54.84 26.87 30.49 29.60
0.20 0.10 24.24 29.03 50.00 34.29 27.85 28.57 0.00 100.00 27.27 35.58 0.00 0.00 55.17 28.81 29.44 27.98
0.20 0.20 22.95 32.18 66.67 30.77 21.74 30.40 0.00 100.00 13.33 38.18 0.00 0.00 62.22 33.33 23.60 24.08
0.20 0.50 13.33 27.03 0.00 22.22 20.69 16.84 0.00 0.00 15.38 41.38 0.00 0.00 51.61 24.56 17.65 20.86
0.50 0.01 21.43 27.74 50.00 46.15 26.67 21.10 0.00 66.67 19.35 34.12 0.00 4.35 53.97 26.09 31.49 29.60
0.50 0.05 22.56 27.94 50.00 42.11 25.29 23.81 0.00 100.00 22.22 34.12 0.00 5.00 54.84 27.07 30.49 29.60
0.50 0.10 25.74 29.27 50.00 33.33 27.85 28.93 0.00 100.00 19.05 35.58 0.00 0.00 55.17 28.33 30.61 28.69
0.50 0.20 23.33 32.18 66.67 30.77 21.74 29.03 0.00 100.00 13.33 38.53 0.00 0.00 62.22 31.17 24.84 24.21
0.50 0.50 9.09 27.03 0.00 23.53 20.69 17.02 0.00 0.00 15.38 41.38 0.00 0.00 51.61 24.56 17.65 20.86
1.00 0.01 21.43 27.74 50.00 46.15 26.67 21.10 0.00 66.67 19.35 34.12 0.00 4.35 53.97 26.09 31.49 29.60
1.00 0.05 22.56 27.94 50.00 42.11 25.29 23.92 0.00 100.00 22.22 34.12 0.00 5.00 54.84 27.07 30.49 29.60
1.00 0.10 26.00 29.27 50.00 33.33 27.85 28.75 0.00 100.00 19.05 35.58 0.00 0.00 54.24 28.33 30.61 28.69
1.00 0.20 23.33 32.56 66.67 30.77 21.74 29.03 0.00 100.00 13.33 38.53 0.00 0.00 62.22 31.17 24.69 24.34
1.00 0.50 9.09 27.03 0.00 23.53 20.69 17.02 0.00 0.00 15.38 41.38 0.00 0.00 51.61 24.56 17.65 20.86
Appendix K Finding the Best In-Context Examples for All Documents
We performed an ablation study to find the global top-$K$ in-context examples for all dev documents, where $K=5$ as in our default REPLM configuration. That is, we search a fixed set of $K$ documents that would yield the best overall performance for a given relation. For this, we leverage the parallel feature selection via group testing (Zhou et al., 2014), where we treat each document in the distantly-supervised set as a feature from the original method and then search for top-$K$ documents.
In essence, this method requires running many experiments, each of which evaluates a random subset of documents, where the number of experiments grows quadratically with respect to the number of documents available. To reduce the number of experiments to a computationally feasible level, we performed this ablation study for only one relation (P118), and we performed our search within the 100 documents that have the highest average cosine similarity to the training documents. This results in 2,659 experiments with different set of in-context examples evaluated on training set.
Table 48 shows the performance of top-$K$ documents selected via group testing (named as REPLM (group testing)). For a better comparison, we include the performance of all methods implemented. For methods with random in-context examples, the performance may be subject to variability across which seed is picked (whereas the performance is deterministic for the other methods), and, hence, we report the standard deviation for this subset of the methods by averaging the performance across 10 runs.
It is shown that the selected documents perform better than the random document selection, e.g., compared to REPLM (random fixed) and REPLM (random all). However, it performs worse than retrieving the most relevant in-context examples for each document, e.g., compared to REPLM (best $\ominus$) and REPLM (best $\oplus$). Therefore, it justifies our original REPLM framework in retrieving the semantically most relevant documents for each dev document.
Table 48: Performance of finding the best in-context examples for all documents via group-testing theory (for relation P118). Shown are F1 scores for each method.
Method F1
REBEL 34.21
REBEL-sent 40.74
REPLM (random fixed) 32.16  $\pm$  5.80
REPLM (random all) 33.51  $\pm$  5.47
REPLM (best $\ominus$) 44.04
REPLM (best $\oplus$) 37.84
REPLM 39.75
REPLM (params adj) 46.28
REPLM (group testing) 36.02
Appendix L Performance with Varying Number of In-context Examples
We performed an ablation study to show how the performance of REPLM framework changes with different number of in-context examples (i. e., varying $K$). For this, we run the same experiment on CONLL04 dataset with $K=3,5,7,9,11$ while keeping all the other parameters fixed. Table 49 shows the performance on each relation type and the overall performance. We observe that more in-context examples yield better F1 score in general. Informed by this observation, we used the highest number of in-context examples ($K$) in our experiments. For the document-level relation extraction dataset DocRED, we use $K=5$ as more than 5 documents do not fit into the context window of our REPLM framework.
Table 49: Comparing the performance across varying number of in-context examples $K$ for the dataset CONLL04. Shown are F1 scores on each relation and overall (Micro) F1 score.
Num. examples ($K$) Kill Live_In Located_In OrgBased_In Work_For Overall
3 88.89 64.04 57.52 67.05 72.85 68.44
5 90.72 63.80 60.36 68.57 78.43 70.54
7 93.75 64.37 56.05 71.66 78.43 70.93
9 90.72 62.96 57.69 71.51 77.03 70.35
11 91.84 64.80 62.96 73.45 80.54 72.94
Appendix M Is REPLM Actually Learning to Extract Relations?
We test whether our REPLM framework is actually learning to extract the relation from the document or simply retrieving the facts from its own memory in the specified format. For this, we re-construct the sentences from CONLL04 (Roth and Yih, 2004) with fake entity names that are not mentioned anywhere in the web. As a result, we push the limits of our REPLM and test whether the model correctly extracts the relations about the entities that appear only in the input sentence. To the best of our knowledge, we are the first to design such an experiment to shed light on the learning abilities of a LM in the context of relation extraction.
Table 50 compares the performance of our REPLM based on the original dataset vs. the same dataset with fake entities. There is only a slight decrease in the overall performance with the fake entities (F1 score of 70.47 vs. 72.94), confirming that our REPLM is actually learning to extract the relations from the context. After this confirmation, we further investigate “how” our REPLM learns to extract relations, by contrasting its behavior against the presence of adversarial in-context examples (see Appendix N).
We conjecture that the slight decrease in the performance can be attributed to mainly two factors: (1) In some cases, the relation between the entities becomes unclear without knowing what these entities actually are. As can be seen from the in-context examples at Sec. M.1, it is hard to infer “work for” relation between “Entity62” and “Entity22” from “ normal-$\ldots$ , said Lt. Entity62 of the Entity22”, which mitigates the model performance. (2) The random entity names (e. g., “Entity10”, “Entity55” etc.) are not seen during the pre-training of the model. Therefore, the likelihood of generating the fake entities from the context is not the same as generating the real names that appeared during the pre-training, which can impact the performance of a LM.
Table 50: Comparing the performance on the original dataset vs. the dataset with random entity names. Shown are F1 scores on each relation and overall (Micro) F1 score.
Dataset Kill Live_In Located_In OrgBased_In Work_For Overall
Original 91.84 64.80 62.96 73.45 80.54 72.94
Random Entities 69.66 76.68 51.53 75.96 75.82 70.47
In the following, we provide an example prompt and the generation of the output based on the re-constructed CONLL04 with the fake entity names.
M.1 Example prompt and generation for the relation Work_For
Input prompt:
Your task is to identify all the unique knowledge triplets of ’Work for’ for a given context. Knowledge triplet will be ordered as relation, subject, and object, which are separated by <==>. If there are multiple triplets, list each of them in a new line. Follow the example context-relation pairs for the formatting of your output.
Context: Entity93 , an Entity79 employee , saw the plane go down as he was getting off work at the airport .
Relation: (Work for <==> Entity93 <==> Entity79)
Context: Entity39 , chief of waste disposal for the Entity97 , said the shafts would be ‘ ‘ the first major holes to be dug into the Entity93 formation. ’ ’
Relation: (Work for <==> Entity39 <==> Entity97)
Context: The fire at the Entity58 Entity97 ’s ammunition dump , which resulted from today ’s explosion , is being localized , according to the fleet ’s senior duty officer Entity55 .
Relation: (Work for <==> Entity55 <==> Entity58)
Context: Entity16 president Entity74 said he did nt know how any criminal wrongdoing could be found in the accident , which developed out of unanticipated conditions , not out of any willful acts.
Relation: (Work for <==> Entity74 <==> Entity16)
Context: The explosion set fire to about Entity55 of wooded area , said Lt. Entity62 of the Entity22 .
Relation: (Work for <==> Entity62 <==> Entity22)
Context: Entity55 ’s Entity10 said he believes the crater was created Entity41 ago when a 50-mile-wide meteorite slammed into the Entity95 .
Output:
Relation: (Work for <==> Entity10 <==> Entity55)
Appendix N Learning Behavior of REPLM with Adversarial In-Context Examples
To further investigate how our REPLM learns from the in-context examples, we evaluate the effect of adversarial in-context examples. To this end, we created the adversarial examples with the following strategy: (1) we randomly selected the half of the documents from the distantly-supervised set, and (2) for the selected documents, we replaced the subject-object pairs of the existing knowledge triplets with random named entities of the same document, which do not share the corresponding relation. We refer to this setup as “50 % clean + 50 % adversarial”. Hence, half of the in-context examples should be adversarial for each dev document. As we do not change the content of the documents, this results in retrieving the same set of in-context examples as in our original framework (i. e., the similarity scores remain the same).
For comparison, we include two other experimental setups: (1) “100 % clean” refers to our original work and (2) “50 % clean” refers to experiments with remaining distantly-supervised documents after discarding the adversarial ones. Of note, we perform all the experiments in this section with our REPLM (best $\ominus$) framework variation. This enables us to directly quantify the impact of the adversarial in-context examples, without any aggregation over the multiple sets of in-context examples.
Table 51 shows the overall results. Furthermore, Tables 52 to 57 reports the performance across each relation type. The performance clearly drops as a result of including the adversarial in-context examples. This has two crucial implications: (1) The quality of the labels is important. Overall, we perform our experiments with distantly-supervised documents that are automatically annotated with certain heuristics to ensure quality (as explained in the main paper). If these are ignored, the resulting in-context examples can mislead the LM. (2) Our framework actually learns from the in-context examples. When it learns the relations with noisy labels, this also reflects into its performance. Further, comparing the performance between “100 % clean” and “50 % clean”, we see that the relation extractions stays at the same level. This informs us that, if some adversarial documents exist in the dataset, filtering them out is a viable option and there is no need to find another method to fix their labels.
Table 51: Document-level relation extraction performance with adversarial in-context examples. Shown: Micro F1 scores.
Data Source F1-Score
100 % clean 31.31
50 % clean 31.16
50 % clean + 50 % adversarial 18.33
Table 52: Performance with adversarial in-context examples. Shown are F1 scores on each relation. (Part 1 / 6)
Data Source P6 P17 P19 P20 P22 P25 P26 P27 P30 P31 P35 P36 P37 P39 P40 P50
100 % clean 35.96 24.60 71.38 50.39 18.18 15.38 33.33 29.14 31.25 6.59 31.46 47.06 31.17 25.00 20.38 38.60
50 % clean 40.91 24.41 72.05 43.17 18.02 20.69 31.94 28.61 39.15 9.20 36.36 50.98 28.21 25.00 18.67 40.24
50 % clean + 50 % adversarial 20.93 11.28 44.91 26.55 8.85 7.41 21.62 16.12 17.73 15.73 9.30 20.00 12.66 14.29 11.25 20.73
Table 53: Performance with adversarial in-context examples. Shown are F1 scores on each relation. (Part 2 / 6)
Data Source P54 P57 P58 P69 P86 P102 P108 P112 P118 P123 P127 P131 P136 P137 P140 P150
100 % clean 48.30 47.62 34.38 57.47 23.26 44.44 30.19 32.65 44.04 30.19 18.49 25.50 20.00 12.90 14.17 31.28
50 % clean 50.77 47.06 33.33 56.98 21.43 41.18 31.25 34.04 42.99 32.38 17.09 23.94 22.22 16.67 9.52 32.14
50 % clean + 50 % adversarial 35.34 25.37 12.50 23.86 11.24 31.14 18.35 17.02 17.39 26.55 17.24 16.12 23.08 13.33 7.52 13.15
Table 54: Performance with adversarial in-context examples. Shown are F1 scores on each relation. (Part 3 / 6)
Data Source P155 P156 P159 P161 P162 P166 P170 P171 P172 P175 P176 P178 P179 P190 P194 P205
100 % clean 23.33 21.51 40.85 33.85 14.12 26.92 10.00 10.53 23.73 40.71 24.14 28.57 22.45 100.00 19.57 13.56
50 % clean 17.24 27.08 37.76 31.89 16.47 30.77 10.67 10.00 29.51 41.77 25.45 30.88 26.26 66.67 15.22 17.86
50 % clean + 50 % adversarial 6.96 16.16 27.78 19.20 16.28 20.18 14.08 17.39 6.78 29.57 14.81 28.79 12.37 50.00 6.59 3.77
Table 55: Performance with adversarial in-context examples. Shown are F1 scores on each relation. (Part 4 / 6)
Data Source P206 P241 P264 P272 P276 P279 P355 P361 P364 P400 P403 P449 P463 P488 P495 P527
100 % clean 14.63 43.24 30.77 35.09 20.00 11.11 30.77 27.12 52.00 40.00 18.52 33.33 38.61 18.18 22.71 23.31
50 % clean 14.52 49.32 34.45 35.71 26.36 10.91 29.63 24.85 34.78 41.90 25.45 35.62 34.20 17.65 22.06 20.19
50 % clean + 50 % adversarial 8.13 25.97 29.03 21.82 9.68 0.00 3.92 15.30 8.16 36.70 18.18 22.54 20.51 12.12 10.13 9.88
Table 56: Performance with adversarial in-context examples. Shown are F1 scores on each relation. (Part 5 / 6)
Data Source P551 P569 P570 P571 P576 P577 P580 P582 P585 P607 P674 P676 P706 P710 P737 P740
100 % clean 36.36 60.95 46.92 38.96 14.49 46.78 27.40 31.11 25.64 26.09 26.42 61.54 16.33 27.96 0.00 38.46
50 % clean 36.36 61.22 47.93 37.58 12.12 43.87 30.56 37.21 23.38 25.41 22.22 46.15 16.33 26.37 0.00 38.46
50 % clean + 50 % adversarial 0.00 33.22 22.38 20.53 5.97 27.52 15.38 18.60 10.39 16.30 16.16 15.38 13.04 8.70 0.00 16.00
Table 57: Performance with adversarial in-context examples. Shown are F1 scores on each relation. (Part 6 / 6)
Data Source P749 P800 P807 P840 P937 P1001 P1056 P1198 P1336 P1344 P1365 P1366 P1376 P1412 P1441 P3373
100 % clean 25.35 31.46 66.67 46.15 21.28 22.06 0.00 100.00 31.58 45.54 0.00 0.00 52.38 42.35 20.00 20.11
50 % clean 19.72 36.78 66.67 46.15 27.91 23.94 0.00 100.00 31.58 42.86 0.00 0.00 58.54 42.86 20.69 24.86
50 % clean + 50 % adversarial 14.71 14.89 0.00 23.08 21.28 8.63 0.00 0.00 9.52 15.38 0.00 0.00 50.00 34.41 13.02 15.96